### v1

# Federated Learning with Reduced Information Leakage and Computation

Abstract
Federated learning (FL) is a distributed learning paradigm that allows multiple decentralized clients to collaboratively learn a common model without sharing local data. Although local data is not exposed directly, privacy concerns nonetheless exist as clients’ sensitive information can be inferred from intermediate computations. Moreover, such information leakage accumulates substantially over time as the same data is repeatedly used during the iterative learning process. As a result, it can be particularly difficult to balance the privacy-accuracy trade-off when designing privacy-preserving FL algorithms. In this paper, we introduce Upcycled-FL, a novel federated learning framework with first-order approximation applied at every even iteration. Under this framework, half of the FL updates incur no information leakage and require much less computation. We first conduct the theoretical analysis on the convergence (rate) of Upcycled-FL, and then apply perturbation mechanisms to preserve privacy. Experiments on real-world data show that Upcycled-FL consistently outperforms existing methods over heterogeneous data, and significantly improves privacy-accuracy trade-off while reducing 48% of the training time on average.
1 Introduction
Federated learning has emerged as an important paradigm for learning models in a distributed fashion, whereby data is distributed across different clients and the goal is to jointly learn a model from the distributed data. This is facilitated by a central server and the model can be learned through iterative interactions between the central server and clients: at each iteration, each client performs certain computation using its local data; the local results are collected and aggregated by the central server; the aggregated result is then sent to clients to update local results; and so on till the learning task is deemed accomplished.
Although each client’s data is not shared directly with the central server, their sensitive information may be inferred from the intermediate computations. Indeed, as the same data is repeatedly used during the iterative learning process, each client’s total information leakage accumulates substantially over time. For randomization-based privacy-preserving techniques such as differential privacy [1], the perturbation required for guaranteeing the privacy of each client also grows as the total information leakage increases. As a result, it can be particularly difficult to balance the privacy and accuracy trade-off in FL.
This paper proposes a novel federated learning framework called Upcycled Federated Learning (Upcycled-FL)***The word “upcycle” refers to reusing material so as to create higher-quality things than the original., in which clients’ information leakage can be reduced such that it only occurs during half of the updates. This is attained by modifying the even iterations of a baseline FL algorithm with first-order approximation, which allows us to compute the update using existing results from previous iterations without using the client’s data. Moreover, the updates in even iterations only involve addition/subtraction operations on existing results at the central server; the computational costs and the training time can be reduced significantly.
To preserve privacy, we can apply existing privacy mechanisms to Upcycled-FL. As an example, we consider differential privacy (DP) because it can (i) rigorously quantify the total privacy leakage for complex algorithms such as FL; (ii) defend against attackers regardless of their background knowledge; (iii) provide heterogeneous privacy guarantees for different clients. Since DP preserves the client’s privacy by randomizing the algorithm (e.g., adding noise to the output, objective, or gradient of local clients), Upcycled-FL by reducing the total information leakage, requires less randomization to attain a certain level of privacy. Therefore, the privacy-accuracy trade-off is also improved.
We emphasize that the idea of upcycling is orthogonal to both the baseline FL algorithm, and the DP perturbation method. In this paper, we adopt FedProx [2] as the baseline FL algorithm and apply first-order approximation to it, but our approach can be applied to other FL frameworks, such as FedNova [3] and pFedMe [4]. For DP perturbation methods, we consider both output and objective perturbation as examples to quantify the privacy loss, while other DP methods can also be used.
It is worth noting that although differentially private federated learning has been extensively studied in the literature, e.g., [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], all these algorithms need client’s local data to update the model and the information leakage inevitably occurs at every iteration. This is fundamentally different from this work, where we propose a novel method for FL that reduces information leakage and computation.
In addition to private federated learning, many approaches were proposed in the DP literature to improve the privacy-accuracy trade-off (e.g., privacy amplification by sampling [21, 22, 23, 24, 25, 26, 27], leveraging non-private public data [28, 29], shuffling [30], using weaker privacy notion [31], using tighter privacy composition analysis tool [27]). However, none of these strategies affect the algorithmic properties of learning algorithms. By contrast, our method improves the privacy-accuracy trade-off by modifying the property of FL algorithm (i.e., reducing the total information leakage); this improvement on the algorithmic property is independent of the privacy notion/mechanism or the analysis method. More related work is discussed in Appendix B
Our main contributions are summarized as follows.
  • We propose Upcycled-FL (Algorithm 2), a novel federated learning framework with reduced information leakage and computation.
  • We conduct convergence (rate) analysis (Section 4, Theorem 4.6) and identify a sufficient condition for the convergence of Upcycled-FL.
  • We apply output perturbation and objective perturbation to Upcycled-FL and conduct privacy analysis (Section 5, Theorem C.1).
  • We evaluate the performance of (private) Upcycled-FL on real-world data (Section 6). The results show that Upcycled-FL, by reducing information leakage, outperforms existing algorithms over heterogeneous data and improves the accuracy-privacy trade-off.
2 Problem Formulation and Preliminaries
Consider an FL system consisting of a central server and a set $\mathcal{I}$ of clients. Each client $i$ has its local data $\mathcal{D}_{i}$ and they can be non-i.i.d across the clients. The goal of FL is to learn a model $\omega\in\mathbb{R}^{d}$ from data $\cup_{i\in\mathcal{I}}\mathcal{D}_{i}$ by solving the following optimization:
$\textstyle\min_{\omega}f(\omega):=\sum_{i\in\mathcal{I}}p_{i}F_{i}(\omega;\mathcal{D}_{i})=\mathbb{E}\left[F_{i}(\omega;\mathcal{D}_{i})\right],$ (1)
where $p_{i}=\frac{|\mathcal{D}_{i}|}{\sum_{j\in\mathcal{I}}|\mathcal{D}_{j}|}$ is the size of client $i$’s data as a fraction of the total data samples, $\mathbb{E}[\cdot]$ is defined as the expectation over clients, $F_{i}(\omega;\mathcal{D}_{i})$ is the local loss function associated with client $i$ and depends on local dataset $\mathcal{D}_{i}$. In this work, we allow $F_{i}(\omega;\mathcal{D}_{i})$ to be possibly non-convex.
FL Algorithm. Let $\omega_{i}^{t}$ be client $i$’s local model parameter at time $t$. In FL, the model is learned through an iterative process: at each time step $t$, (1) local computations: each active client updates its local model $\omega_{i}^{t}$ using its local data $\mathcal{D}_{i}$; (2) local models broadcasts: local models (or gradients) are then uploaded to the central server; (3) model aggregation: the central server aggregates results received from clients to update the global model parameter $\overline{\omega}^{t}=\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t}$; (4) model updating: the aggregated model is sent back to clients and is used for updating local models at $t+1$. In this work, we consider a baseline FL framework FedProx and apply the proposed method to it, while our method can also be used on other FL frameworks such as FedNova [3] and pFedMe [4]. The details of FedProx are given in Algorithm 1.
Algorithm 1 FedProx [2]
1:  Input: $\mu>0$, $\{\mathcal{D}_{i}\}_{i\in\mathcal{I}}$, $\overline{\omega}^{0}$
2:  for $t=1$ to $T$ do
3:     The central server sends the current global model parameter $\overline{\omega}^{t}$ to all the clients.
4:     A subset of clients get active and each active client updates its local model by finding (approximate) minimizer of local loss function:
$\omega_{i}^{t+1}=\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{t}||^{2}.$
5:     Each client sends its local model to server.
6:     The central server updates the global model by aggregating all local models:
$\overline{\omega}^{t+1}=\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t+1}.$
7:  end for
During the learning process, each client’s local computation is exposed to third parties at every iteration: its models/gradients need to be uploaded to the central server, and the global models calculated based on them are shared with all clients. It is thus critical to ensure the FL is privacy-preserving. In this work, we consider differential privacy as the notion of privacy.
Differential Privacy [1]. Differential privacy (DP) centers around the idea that the output of a certain computational procedure should be statistically similar given singular changes to the input, thereby preventing meaningful inference from observing the output.
In FL, the information exposed by each client $i$ includes all intermediate computations $\{\omega_{i}^{t}\}_{t=1}^{T}$. Consider a randomized FL algorithm $\mathcal{A}(\cdot)$ that generates a sequence of private local models $\{\widehat{\omega}_{i}^{t}\}_{t=1}^{T}$, we say it satisfies DP for client $i$ over $T$ iterations if the following holds for any possible output $O\in\mathbb{R}^{d}\times\cdots\times\mathbb{R}^{d}$, and for any two neighboring local datasets $\mathcal{D}_{i}$, $\mathcal{D}^{\prime}_{i}$:
$\Pr(\{\widehat{\omega}_{i}^{t}\}_{t=0}^{T}\in O|\mathcal{D}_{i})\leq\exp{(\varepsilon)}\cdot\Pr(\{\widehat{\omega}_{i}^{t}\}_{t=0}^{T}\in O|\mathcal{D}^{\prime}_{i})+\delta.$
where $\varepsilon\in[0,\infty)$ bounds the privacy loss, and $\delta\in[0,1]$ loosely corresponds to the probability that algorithm fails to bound the privacy loss by $\varepsilon$. Two datasets are neighboring datasets if they are different in at most one data point.
3 Proposed Algorithm: Upcycled-FL
Main idea: Fundamentally, the accumulation of information leakage over iterations stems from the fact that the client’s data $\mathcal{D}_{i}$ is used in every update. If the updates can be made without directly using this original data, but only from computational results that already exist, then the information leakage originating from these updates will be zero, and meanwhile the computational cost may be reduced significantly. Based on this idea, we propose Upcycled-FL, which modifies FedProx such that the earlier computations are repeatedly used for new updates.
Upcycling Information: Next, we present Upcycled-FL and illustrate how the client’s total information leakage is reduced under this method.
We modify FedProx by applying first-order approximation to even iterations. Specifically, at even iteration $2m$, we expand $F_{i}(\omega;\mathcal{D}_{i})$ at $\omega^{2m-1}_{i}$ (local model in the previous iteration). Based on Taylor series expansion, we have:
$F_{i}(\omega;\mathcal{D}_{i})$ $=$ $F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})$
$+\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}(\omega-\omega^{2m-1}_{i})$
$+\mathcal{O}(||\omega-\omega^{2m-1}_{i}||^{2}).$
We approximate term $\mathcal{O}(||\omega-\omega^{2m-1}_{i}||^{2})$ by $\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}$ for some constant $\lambda_{m}\geq 0$, which can be different for different iteration $2m$, then the even update of FedProx under such an approximation becomes:
$\omega_{i}^{2m}$ $=$ $\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2}$ (2)
$\approx$ $\arg\min_{\omega}F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})+\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}\omega$
$+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2}$
$=$ $\arg\min_{\omega}\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}\omega$
$+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2}.$
Note that in the above even update, the only term that depends on dataset $\mathcal{D}_{i}$ is $\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})$, which can be derived directly from the previous odd iteration. Specifically, according to the first-order condition, the following holds at odd iterations:
$\omega_{i}^{2m-1}=\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-2}||^{2}$
[…]
Plug $\nabla F_{i}(\omega_{i}^{2m-1};\mathcal{D}_{i})$ into even updates (2), we have
$\omega_{i}^{2m}$$\approx$ $\arg\min_{\omega}\mu(\overline{\omega}^{2m-2}-\omega_{i}^{2m-1})^{T}\omega$ (4)
$+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2}.$
By first-order condition, even updates (4) can be reduced to:
$\omega_{i}^{2m}\approx\omega^{2m-1}_{i}+\frac{\mu}{\mu+\lambda_{m}}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right).$ (5)
It turns out that with first-order approximation, dataset $\mathcal{D}_{i}$ is not used in the even updates. Furthermore, these updates can be conducted at the central server directly. That is, the central server updates the global model by aggregating:
$\overline{\omega}^{2m}$ $=$ $\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{2m}$
$\approx$ $\sum_{i\in\mathcal{I}}p_{i}\omega^{2m-1}_{i}+\frac{\mu}{\mu+\lambda_{m}}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right)$
$=$ $\overline{\omega}^{2m-1}+\frac{\mu}{\mu+\lambda_{m}}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right)$
Therefore, even updates only involve addition/subtraction operations on the existing global models from previous iterations (i.e., $\overline{\omega}^{2m-1},\overline{\omega}^{2m-2}$): the computational cost and training time is reduced significantly. Note that the first-order approximation is only applied to even iterations, while the odd iterations should remain the same as FedProx to ensure Eqn. (3) holds. The entire updating procedure of Upcycled-FL is summarized in Algorithm 2.
Because $\mathcal{D}_{i}$ is only used in odd iterations, information leakage only happens during odd updates. Intuitively, the reduced information leakage would require less perturbation to attain a certain level of privacy guarantee, which further results in higher accuracy and improved privacy-accuracy trade-off. In the following sections, we first analyze the convergence property of Upcycled-FL and then apply privacy mechanisms to satisfy differential privacy.
Discussion: Indeed, if we view two consecutive steps (odd $12m-$ 1 followed by even $2m$) of Upcycled-FL as a single iteration $t$, then Upcycled-FL and FedProx will incur the same information leakage but will differ at the phase of global model aggregation. Specifically, instead of simply aggregating the global model by averaging over local updates (i.e., $\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t}$) as in FedProx, Upcycled-FL not only takes the average of local updates but also push the aggregation moving more toward the updating direction (i.e., $\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t}-\overline{\omega}^{t-1}$). In Figure 1, we illustrate the difference between FedProx and Upcycled-FL from the two perspectives: (i) Upcycled-FL can be regarded as reusing the intermediate updates of FedProx, compared to which the total information leakage is reduced; or (ii) Upcycled-FL can be regarded as FedProx with a different global aggregation rule, and the larger global update in Upcycled-FL accelerates the learning process with the same information leakage.
(a) Upcycled-FL reuses the intermediate updates
(b) Upcycled-FL uses different aggregation rule
Figure 1: Consider Upcycled-FL from two perspectives
4 Convergence Analysis
In this section, we analyze the convergence property of Upcycled-FL. Note that we do not require local functions $F_{i}(\cdot)$ to be convex. Moreover, we consider practical settings where data are non-i.i.d across different clients. Similar to [2], we introduce a measure below to quantify the dissimilarity between clients in the federated network.
Definition 4.1 ($B$-Dissimilarity [2]).
The local loss function $F_{i}$ is $B$-dissimilar if $\forall\omega$, we have $\mathbb{E}[||\nabla F_{i}(\omega)||^{2}]\leq||\nabla f(\omega)||^{2}B^{2}$.
where $\mathbb{E}[\cdot]$ denotes the expectation over clients (Eqn. (1)). Parameter $B\geq 1$ captures the statistical heterogeneity across different clients: when all clients are homogeneous with i.i.d. data, we have $B=1$ for all local functions; the larger value of $B$, the more dissimilarity among clients.
Assumption 4.2.
Local loss functions $F_{i}$ are $B$-dissimilar and $L$-Lipschitz smooth.
Note that $B$-dissimilarity can be satisfied if the divergence between the gradient of the local loss function and that of the aggregated global function is bounded, as stated below.
Lemma 4.3.
$\forall i$, there exists $B$ such that $F_{i}$ is $B$-dissimilar if $||\nabla F_{i}(\omega)-\nabla f(\omega)||\leq\kappa_{i},\forall\omega$ for some $\kappa_{i}$.
Assumption 4.4.
$\forall i$, $h_{i}(\omega;\overline{\omega}^{t}):=F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{t}||^{2}$ are $\rho$-strongly convex.
The above assumptions are fairly standard. They first appeared in [2] and are adopted in subsequent works such as [4, 32, 33].
Note that strongly convex assumption is not on local objective $F_{i}(\omega;\mathcal{D}_{i})$, but the regularized function $F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{t}||^{2}$, i.e., the assumption can be satisfied by selecting a sufficiently large $\mu>0$. Indeed, as shown in Section 6, our algorithm still converges even when Assumption 4.2 and 4.4 do not hold (e.g., DNN). Next, we conduct the theoretical analysis on the convergence (rate) of Upcycled-FL.
Algorithm 2 Proposed framework: Upcycled-FL
1:  Input: $\lambda_{m}>0$, $\mu>0$, $\{\mathcal{D}_{i}\}_{i\in\mathcal{I}}$, $\overline{\omega}^{0}$
2:  for $m=1$ to $M$ do
3:     The central server sends the current global model parameter $\overline{\omega}^{2m-2}$ to all the clients.
4:     A subset of clients are selected to be active and each active client updates its local model by finding (approximate) minimizer of local loss function:
$\omega_{i}^{2m-1}\leftarrow\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-2}||^{2}.$
5:     Clients send local models to the central server.
6:     The central server updates the global model by aggregating all local models:
$\overline{\omega}^{2m-1}$ $=$ $\textstyle\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{2m-1}$
$\overline{\omega}^{2m}$ $=$ $\overline{\omega}^{2m-1}+\frac{\mu}{\mu+\lambda_{m}}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right)$
7:  end for
Lemma 4.5.
Let $\mathcal{S}_{m}$ be a set of $K$ randomly selected clients got updated (i.e., active clients) at iterations $12m-$ 1 and $2m$, and $\mathbb{E}_{\mathcal{S}_{m}}[\cdot]$ be the expectation with respect to the choice of clients. Then under Assumption 4.2 and 4.4, we have
$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]$$\leq$ $f(\overline{\omega}^{2m-1})-\boldsymbol{C_{1}}||\nabla f(\overline{\omega}^{2m-1})||^{2}$
$+\boldsymbol{C_{2}}\cdot h^{1}_{m}+\boldsymbol{C_{3}}\cdot h^{2}_{m},$
where
$h^{1}_{m}$ ::=:= $||\nabla f(\overline{\omega}^{2m-1})||\cdot||\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}||;$
$h^{2}_{m}$ ::=:= $||\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}||^{2}.$
The details of the term $\boldsymbol{C_{1}}$, $\boldsymbol{C_{2}}$, $\boldsymbol{C_{3}}$ (expressed as functions of $L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{1}{K},\frac{\mu}{\lambda_{m}}$) are in Appendix E, Eqn. (10)-(12).
Lemma 4.5 characterizes the relation of the values of global objective function over two consecutive odd iterations. It is easy to verify $\boldsymbol{C_{2}},\boldsymbol{C_{3}}\geq 0$. By rearranging and telescoping, we get the following convergence rate of Upcycled-FL.
Theorem 4.6 (Convergence rate of Upcycled-FL).
Under Assumption 4.2 and 4.4, if $\boldsymbol{C_{1}}>0$, we have
$\min_{m\in[M]}\mathbb{E}[||\nabla f(\overline{\omega}^{2m-1})||^{2}]$
$\leq$ $\frac{1}{M}\sum_{m=0}^{M}\mathbb{E}[||\nabla f(\overline{\omega}^{2m-1})||^{2}]$
$\leq$ $\frac{f(\overline{\omega}^{0})-f(\overline{\omega}^{*})}{M\boldsymbol{C_{1}}}+\frac{\sum_{m=0}^{M}\boldsymbol{C_{2}}h^{1}_{m}}{M\boldsymbol{C_{1}}}+\frac{\sum_{m=0}^{M}\boldsymbol{C_{3}}h^{2}_{m}}{M\boldsymbol{C_{1}}},$
where $\overline{\omega}^{0}$ and $\overline{\omega}^{*}$ denote the initial and the optimal global model parameters, and $\boldsymbol{C_{2}}$ and $\boldsymbol{C_{3}}$ are decreasing in $\frac{\lambda_{m}}{\mu}$.
Theorem 4.6 implies that tunable $\mu$ and $\lambda_{m}$ are key hyper-parameters that control the convergence (rate) and robustness of Upcycled-FL. Recall that $\mu$ penalizes the deviation of local model $\omega_{i}^{2m}$ from global aggregated model $\overline{\omega}^{2m-1}$, while $\lambda_{m}$ penalizes the update of local model $\omega_{i}^{2m}$ from its previous update $\omega_{i}^{2m-1}$.
Because $\boldsymbol{C_{1}}:=C_{1}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{1}{K}\Big{)}$ does not depend on $\lambda_{m}$ (by Eqn. (10)), for proper $\mu$ and local functions $F_{i}$, the condition $\boldsymbol{C_{1}}>0$ in Theorem 4.6 can hold for any $\lambda_{m}$. However, $\lambda_{m}$ could affect the convergence rate via terms $\boldsymbol{C_{2}}:=C_{2}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{1}{K},\frac{\mu}{\mu+\lambda_{m}}\Big{)}$ and $\boldsymbol{C_{3}}:=C_{3}\Big{(}L,\frac{1}{\mu},\frac{1}{\rho},\frac{1}{K},\frac{\mu}{\mu+\lambda_{m}}\Big{)}$. Specifically, as the ratio $\frac{\lambda_{m}}{\mu}$ increases, both $\boldsymbol{C_{2}}$ and $\boldsymbol{C_{3}}$ decrease (by Eqn. (11)-(12)) which results in a tighter convergence rate bound. We empirically examine the impacts of $\mu$ and $\lambda_{m}$ in Section 6.
It is worth noting that the convergence rate also depends on data heterogeneity, captured by dissimilarity $B$. According to Eqn. (10), $\boldsymbol{C_{1}}>0$ must hold when $B=0$ (i.i.d. clients). Although $\boldsymbol{C_{1}}$ may become negative as $B$ increases, the experiments in Section 6 show that Upcycled-FL can still converge when data is highly heterogeneous.
Assumption 4.7.
$||\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}||\leq h,\forall m$ and $||\nabla f(\omega)||\leq d,\forall\omega$.
Assumption 4.7 is standard and has been used when proving the convergence of FL algorithms [34, 35]; it requires that the difference of aggregated weights between two consecutive iterations and the gradient $||\nabla f(\omega)||$ are bounded. Under this assumption, we have the following corollary.
Corollary 4.8 (Convergence to the stationary point).
Under Assumption 4.2, 4.4, and 4.7, for fixed $\mu,K$, if $\lambda_{m}$ is taken such that $\frac{\mu}{\mu+\lambda_{m}}=\mathcal{O}\Big{(}\frac{1}{\sqrt{M}}\Big{)}$, then the convergence rate of Upcycled-FL reduces to $\mathcal{O}(\frac{1}{\sqrt{M}})$.
Corollary 4.8 provides guidance on how to choose the value of $\lambda_{m}$ to guarantee the convergence of Upcycled-FL, i.e., by taking an increasing sequence of $\{\lambda_{m}\}_{m=1}^{M}$. Intuitively, increasing $\lambda_{m}$ during the training helps stabilize the algorithm, because the deviation of local models from the previous update gets penalized more under a larger $\lambda_{m}$.
5 Private Upcycled-FL
In this section, we present a privacy-preserving version of Upcycled-FL. Many perturbation mechanisms can be adopted to achieve differential privacy such as objective perturbation [36, 37], output perturbation [36, 38], gradient perturbation [39, 40], etc. Next, we will use output and objective perturbation as examples to illustrate that Upcycled-FL, by reducing the total information leakage, has a better privacy-accuracy trade-off than other methods such as FedProx, FedAvg, etc. Note that both output and objective perturbation methods are used for generating private updates at odd iterations, which can be used directly for even updates.
Output perturbation: the private odd updates $\widehat{\omega}_{i}^{2m-1}$ are generated by first clipping the local models ${\omega}_{i}^{2m-1}$ and then adding a noise random vector $n_{i}^{m}$ to the clipped model:
Clip odd update: $\xi({\omega}_{i}^{2m-1})=\dfrac{{\omega}_{i}^{2m-1}}{\max\left(1,\frac{||{\omega}_{i}^{2m-1}||_{2}}{\tau}\right)}$
Perturb with noise: $\widehat{\omega}_{i}^{2m-1}=\xi({\omega}_{i}^{2m-1})+n_{i}^{m}$
where parameter $\tau>0$ is the clipping threshold; the clipping ensures that if $||{\omega}_{i}^{2m-1}||_{2}\leq\tau$, then update remains the same, otherwise it is scaled to be of norm $\tau$.
Objective perturbation: a random linear term $\langle n^{m}_{i},\omega\rangle$ is added to the objective function in odd $(2m+1)$-th iteration, and the private local model $\widehat{\omega}_{i}^{2m+1}$ is found by solving a perturbed optimization:
$\widehat{\omega}_{i}^{2m+1}=\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{2m}||^{2}+\langle n^{m}_{i},\omega\rangle,$
Given noisy $\widehat{\omega}_{i}^{2m-1}$ generated by either method, the private even updates $\widehat{\omega}_{i}^{2m}$ can be computed directly at the central server using the noisy aggregation $\sum_{i\in\mathcal{I}}p_{i}\widehat{\omega}_{i}^{2m-1}$.
Privacy Analysis. Next, we conduct privacy analysis and theoretically quantify the total privacy loss of private Upcycled-FL. Because even updates are computed directly using already private intermediate results without using dataset $\mathcal{D}_{i}$, no privacy leakage occurs at even iterations. This can be formally stated as the following lemma.
Lemma 5.1.
For any $m=1,2,\cdots$, if the total privacy loss up to $12m-$ 1 can be bounded by $\varepsilon_{m}$, then the total privacy loss up to the $2m$-th iteration can also be bounded by $\varepsilon_{m}$.
Lemma 5.1 is straightforward; it can be derived directly by leveraging a property of differential privacy called immunity to post-processing [41], i.e., a differentially private output followed by any data-independent computation remains satisfying differential privacy.
Based on Lemma 5.1, we can quantify the total privacy loss of private Upcycled-FL. We shall adopt moments accountant method [27] for output perturbation, and the analysis method in [36, 42, 43] for objective perturbation. These theoretical results are presented in Appendix C, Theorem C.1-C.2.
6 Experiments
In this section, we empirically evaluate Upcycled-FL and compare it with two algorithms: FedAvg [44] and FedProx [2]. We first consider non-private algorithms to examine the convergence (rate) and robustness of Upcycled-FL against statistical/device heterogeneity. Then, we adopt both output and objective perturbation to evaluate the private Upcycled-FL.
Experimental setup. We run experiments on both synthetic and real data.
  • Synthetic data: using the method in [2], we generate Synthetic(iid), Synthetic(0,0), Synthetic(0.5,0.5), Synthetic(1,1), four synthetic datasets with increasing statistical heterogeneity.
  • Real data: we adopt two real datasets: 1) Femnist, a federated version of Emnist [45]; 2) Sentiment140(Sent140), a text sentiment analysis task on tweets [46]. Both of the datasets can be obtained from LEAF [47].
To simulate device heterogeneity, we randomly select a fraction of devices to train at each round, and assume there are stragglers that cannot train for full rounds; both devices and stragglers are selected by random seed to ensure they are the same for all algorithms. We also consider several learning algorithms: we learn logistic regression on synthetic data; multilayer perceptron (MLP) and deep CNN on Femnist; Stacked LSTM on Sent140. For Femnist, we present the results on MLP here and other results in Appendix H.
Because even iterations of Upcycled-FL only involve addition/subtraction operations with almost no computational cost, we train Upcycled-FL with double iterations compared to FedAvg and FedProx in approximately same training time. Unless explicitly stated, the results we report are averaged outcomes over all devices. More details of experimental setup are in Appendix H.1.
(a) Synthetic(0.5,0.5)
(b) Femnist
Figure 2: Convergence of Upcycled-FL, FedProx, FedAvg under the approximate same training time (left) and impact of $\lambda_{m}$ on convergence rate of Upcycled-FL (right).
Table 1: Training(testing) accuracy of FedAvg, FedProx, Upcycled-FL with 90% straggler: FedAvg/FedProx/Upcycled-FL are trained over synthetic data for 80/80/160 iterations, and are trained over Femnist for 150/150/300 iterations. For Upcycled-FL, $\lambda=0.04,0.12,0.21,0.43$ is set based on the value of $\mu=0.1,0.3,0.5,1$, respectively. Upcycled-FL significantly outperforms FedAvg and FedProx.
Dataset Proximal term Accuracy(%)
FedAvg FedProx Upcycled-FL
Synthetic(iid) $\mu=0.1$ 97.80%(97.51%) 96.86%(96.34%) 97.88%(97.51%)
$\mu=0.3$ 95.18%(94.88%) 96.31%(96.05%)
$\mu=0.5$ 94.03%(93.70%) 95.40%(95.02%)
$\mu=1$ 92.14%(91.80%) 93.76%(93.70%)
Synthetic(0,0) $\mu=0.1$ 78.41%(76.27%) 81.66%(79.50%) 84.41%(82.44%)
$\mu=0.3$ 81.73%(79.35%) 83.83%(81.58%)
$\mu=0.5$ 81.55%(79.35%) 82.84%(81.29%)
$\mu=1$ 81.22%(78.85%) 82.41%(80.57%)
Synthetic(0.5,0.5) $\mu=0.1$ 79.53%(78.97%) 79.63%(79.58%) 81.19%(81.54%)
$\mu=0.3$ 80.31%(80.56%) 82.14%(83.74%)
$\mu=0.5$ 79.55%(79.46%) 82.61%(82.64%)
$\mu=1$ 80.11%(80.20%) 82.74%(82.64%)
Synthetic(1,1) $\mu=0.1$ 69.70%(71.12%) 79.20%(78.83%) 77.38%(78.64%)
$\mu=0.3$ 74.77%(75.91%) 80.43%(81.47%)
$\mu=0.5$ 75.36%(76.67%) 82.78%(83.07%)
$\mu=1$ 75.69%(76.58%) 82.88%(82.69%)
Femnist $\mu=0.1$ 20.43%(20.01%) 82.86%(83.73%) 86.37%(85.91%)
$\mu=0.3$ 76.10%(76.66%) 81.94%(82.12%)
$\mu=0.5$ 68.28%(68.87%) 77.77%(77.60%)
$\mu=1$ 28.82%(29.05%) 54.81%(54.00%)
Sent140 $\mu=0.1$ 73.41%(71.13%) 75.29%(72.57%) 78.12%(72.38%)
$\mu=0.3$ 72.86%(71.38%) 74.27%(73.02%)
$\mu=0.5$ 70.18%(69.07%) 72.92%(71.44%)
$\mu=1$ 62.10%(62.13%) 63.91%(64.11%)
Convergence and Heterogeneity. We evaluate the convergence rate, training time, and accuracy of Upcycled-FL under different parameter settings. We take $\lambda_{m}=\lambda,\forall m$ in experiments, while we observe similar results for time-varying $\lambda_{m}$. Table 1 compares the training/testing accuracy of different algorithms when the device heterogeneity is high (90% stragglers). The results show that Upcycled-FL outperforms both FedProx and FedAvg on all datasets (with appropriate $\mu$). While FedAvg achieves good performance on Synthetic(iid), it is not robust to heterogeneous data. When data is i.i.d., adding the proximal term $\frac{\mu}{2}||\omega-\overline{\omega}^{t}||^{2}$ may hurt the performance. However, the proximal term can help stabilize the algorithm and significantly improve the performance in practical settings when data is heterogeneous; these observations are consistent with [2]. Importantly, Upcycled-FL is more robust to the level of statistical heterogeneity than FedProx as it could attain consistent improvements for all settings. In practice, $\mu$ needs to be carefully tuned; with a carefully selected $\mu$, Upcycled-FL could achieve $2\%\sim 7\%$ improvement on accuracy compared to FedProx.
Figure 2(a) and 2(b) illustrate the convergence property of Upcycled-FL on Synthetic(0.5,0.5) and Femnist (in comparison with FedProx and FedAvg), and examine the impact of hyper-parameter $\lambda_{m}$. Loss in $y$-axes indicates the averaged loss over all devices. In each iteration, 30% of devices are selected with 90% stragglers. We observe that Upcycled-FL could attain a comparable performance as FedProx in the first 80 (resp. 240) iterations on Synthetic(0.5,0.5) (resp. Femnist), with 50% (resp. 20%) of SGD computation reduced. Specifically, the final training (testing) accuracy attained on Synthetic(0.5,0.5) is 82.61% (82.64%) for Upcycled-FL, 79.55% (79.46%) for FedProx, and 79.53% (78.97%) for FedAvg; while the accuracy attained on Femnist is 81.94% (82.12%) for Upcycled-FL, 76.10% (76.66%) for FedProx, and 20.43% (20.01%) for FedAvg. More results on other datasets are presented in Appendix H.4.
(a) Synthetic(iid)
(b) Synthetic(0,0)
(c) Synthetic(iid)
(d) Synthetic(0,0)
Figure 3: Comparison of private Upcycled-FL and private FedProx using objective perturbation ((a)(b)) and output perturbation ((c)(d)) with $\delta=0.001$. We take $\mu=0.1,\lambda=0.04$ for all datasets. $\bar{\epsilon}$ denotes the average $\epsilon$ among all devices. In both experiments, the privacy of private Upcycled-FL is strictly stronger than private FedProx.
The effect of hyper-parameter $\lambda_{m}=\lambda$ on the convergence rate of Upcycled-FL is also evaluated in Figure 2(a)2(b) (right). Although the sufficient conditions in Theorem 4.6 and Corollary 4.8 suggest that Upcycled-FL converges when $\lambda$ is sufficiently large, the experiments indicate that Upcycled-FL can easily converge in practice even with small $\lambda$. Furthermore, we observe that models with smaller $\lambda$ tend to converge faster. This is because that $\lambda$ controls step size $\frac{\mu}{\mu+\lambda}$ (by Eqn. (5)): a larger $\lambda$ results in a smaller step size and thus slower convergence.
Privacy-Accuracy Tradeoff. We next inspect the accuracy-privacy trade-off of private Upcycled-FL and compare it with private FedProx. We adopt both objective and output perturbation to preserve privacy in the experiments, while other privacy-preserving techniques can also be used. For each parameter setting, we perform 10 independent runs of the experiment and record both the mean and range of their losses. To precisely quantify privacy, we focus on settings without device heterogeneity (no straggler) and low statistical heterogeneity (Synthetic(iid), Synthetic(0,0)).
Figure 3(a) and 3(b) demonstrate the performance (averaged loss and the respective privacy loss $\epsilon$) of (non-)private Upcylced-FL and private FedProx on synthetic data using objective perturbation. As expected, private Upcycled-FL is much more stable and significantly improves both privacy and accuracy over private FedProx. This is because the private Upcycled-FL with less information leakage requires much less perturbation to attain the same privacy guarantee as private FedProx. In general, we observe that private Upcycled-FL can experience 9.5% less privacy loss than private FedProx, and meanwhile having 11.0% (or 11.5%) higher training (or testing) accuracy on average and reducing 48% of training time. We also conduct similar experiments using output perturbation (Figure 3(c) and 3(d)). Results show that on average, private Upcycled-FL can experience 12.0% less privacy loss than private FedProx while having 6.1% (or 6.2%) higher training (or testing) accuracy; the training time is still reduced by 48% under Upcycled-FL.
We also conduct experiments to examine the accuracy-privacy trade-off under different magnitudes of perturbation, the results are presented in Appendix H.5.1. It shows that private Upcycled-FL can consistently outperform private FedProx under different levels of privacy.
7 Conclusion
This paper proposes Upcycled-FL, a novel federated learning framework under which information leakage and computation cost can be reduced significantly. We theoretically examined the convergence (rate) of Upcycled-FL. The experiments show that compared to existing algorithms, Upcycled-FL can be more robust on heterogeneous data and attain much better privacy-accuracy trade-off under common differential privacy mechanisms.
Appendix A Notation Table
$\mathcal{I}$ set of agents
$\mathcal{D}_{i}$ dataset of agent $i$
$p_{i}$ size of agent $i$’s data as a fraction of total data samples
$\omega_{i}^{t}$ agent $i$’s local model parameter at time $t$
$\overline{\omega}^{t}$ aggregated model at central server at $t$
$\widehat{\omega}^{t}_{i}$ differentially private version of $\omega_{i}^{t}$
$F_{i}$ local objective function of agent $i$
$f$ overall objective function
$n_{i}^{t}$ random noise added to agent $i$ at time $t$
$\mu$ hyper-parameter for proximal term in FedProx and Upcycled-FL
$\lambda_{m}$ hyper-parameter for first-order approximation at even iteration $2m$ in Upcycled-FL
Appendix B Related Work
Differentially private federated learning. Differential privacy has been widely used in federated learning to provide privacy guarantees [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. For example, [7] uses the Gaussian mechanism for a federated learning problem and propose an incentive mechanism to encourage users to share their data and participate in the training process. [8] introduces $f$-differential privacy, a generalized version of Gaussian differential privacy, and propose a federated learning algorithm satisfying this new notion. [9] proposes a new mechanism called Random Response with Priori (RRP) to achieve local differential privacy and apply this mechanism to the text data by training a Latent Dirichlet Allocation (LDA) model using a federated learning algorithm. [20] adapts the Bayesian privacy accounting method to the federated setting and propose joint accounting method for estimating client-level and instance-level privacy simultaneously and securely. [19] presents a private scheme that adds noise to parameters at the random selected devices before aggregating and provides a convergence bound. [10] combines the Gaussian mechanism with gradient clipping in federated learning to improve the privacy-accuracy tradeoff. [5] considers a different setting where only the last update is publicly released and the central server and other devices are assumed to be trustworthy. However, all these algorithms need client’s local data to update the model and the information leakage inevitably occurs at every iteration. This is fundamentally different from Upcycled-FL, which reuses existing results to update half of iterations and significantly reduces information leakage and computation.
Tackling heterogeneity in federated learning. It’s worth mentioning that, empirically Upcycled-FL also outperforms existing baseline algorithm under device and statistical heterogeneity. In real-world scenarios, local data are often non-identically distributed across different devices; different devices are also often equipped with different specifications and computation capabilities. Such heterogeneity often causes instability in the model performance and leads to divergence. Many approaches have been proposed to tackle this issue. For example, FedAvg [44] uses a random selection of devices at each iteration to reduce the negative impact of statistical heterogeneity; however, it may fail to converge when heterogeneity increases. Other methods include FedProx [2], a generalization and re-parameterization of FedAvg that adds a proximal term to the objective function to penalize deviations in the local model from the previous aggregation, and FedNova [3] that re-normalizes local updates before updating to eliminate objective inconsistency. It turns out that Upcycled-FL exhibits superior performance in the presence of heterogeneity because gradients encapsulate information on data heterogeneity, the reusing of which leads to a boost in performance.
Appendix C Privacy analysis and results
We provide the privacy analysis of private Upcycled-FL under output and objective perturbation.
Here, we focus on settings where local loss function $F_{i}(\omega_{i},\mathcal{D}_{i}):=\frac{1}{|\mathcal{D}_{i}|}\sum_{d\in\mathcal{D}_{i}}\hat{F}_{i}(\omega_{i},d)$ for some $\hat{F}_{i}$, and the guarantee of privacy is presented in Theorem C.1 (output perturbation) and C.2 (objective perturbation) below. The total privacy loss in the following theorem is composed using moments accountant method [27].
Theorem C.1.
Consider the private Upcycled-FL over $2M$ iterations under output perturbation with noise $n_{i}^{m}\sim\mathcal{N}(0,\sigma^{2}\textbf{I})$, then for any $\varepsilon\geq\frac{M\tau^{2}}{2\sigma^{2}|\mathcal{D}_{i}|^{2}}$, the algorithm is $(\varepsilon,\delta)$-DP for agent $i$ for
$\delta=\exp\left(-\frac{M\tau^{2}}{2\sigma^{2}|\mathcal{D}_{i}|^{2}}\Big{(}\frac{\varepsilon\sigma^{2}|\mathcal{D}_{i}|^{2}}{M\tau^{2}}-\frac{1}{2}\Big{)}^{2}\right).$
Equivalently, for any $\delta\in[0,1]$, the algorithm is $(\varepsilon,\delta)$-DP for agent $i$ for
$\varepsilon=2\sqrt{\frac{M\tau^{2}}{2\sigma^{2}|\mathcal{D}_{i}|^{2}}\log(\frac{1}{\delta})}+\frac{M\tau^{2}}{2\sigma^{2}|\mathcal{D}_{i}|^{2}}.$
Theorem C.2.
Consider the private Upcycled-FL over $2M$ iterations under objective perturbation with noise $n^{m}_{i}\sim\exp{(-\alpha_{i}^{m}||n_{i}^{m}||_{2})}$. Suppose $\hat{F}_{i}$ is generalized linear model [48, 39]In supervised learning, the sample $d=(x,y)$ corresponds to the feature and label pair. Function $\hat{F}_{i}(\omega,d)$ is generalized linear model if it can be written as a function of $\omega^{T}x$ and $y$. that satisfies $||\nabla\hat{F}_{i}(\omega;d)||<u_{1},~{}|\hat{F}^{\prime\prime}_{i}|\leq u_{2}$. Let feature vectors be normalized such that its norm is no greater than 1, and suppose $u_{2}\leq 0.5|\mathcal{D}_{i}|\mu$ holds. Then the algorithm satisfies $(\varepsilon,0)$-DP for agent $i$ where $\varepsilon=\sum_{m=0}^{M}\frac{2\alpha^{m}_{i}u_{1}\mu+2.8u_{2}}{|\mathcal{D}_{i}|\mu}.$
The assumptions on $\hat{F}_{i}$ are again fairly standard in the literature, see e.g., [36, 42, 43]. Theorem C.1 and C.2 show that the total privacy loss experienced by each agent accumulates over iterations and privacy loss only comes from odd iterations. In contrast, if consider differentially private FedProx, accumulated privacy loss would come from all iterations. Therefore, to achieve the same privacy guarantee, private Upcycled-FL requires much less perturbation per iteration than private FedProx. As a result, accuracy can be improved significantly. Experiments in Section 6 show that Upcycled-FL significantly improves privacy-accuracy trade-off compared to other methods.
Appendix D Lemmas
Lemma D.1.
Define $\widetilde{\omega}^{t}=\mathbb{E}_{i}(\omega_{i}^{t})$. Suppose conditions in Theorem 4.6 hold. Then the following holds for all $m$:
$f(\widetilde{\omega}^{2m+1})$$\leq$ $f(\overline{\omega}^{2m-1})-\widehat{C}_{1}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho}\Big{)}||\nabla f(\overline{\omega}^{2m-1})||^{2}$
$+\widehat{C}_{2}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{\mu}{\mu+\lambda_{m}}\Big{)}||\nabla f(\overline{\omega}^{2m-1})||\cdot||\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}||$
$+\widehat{C}_{3}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{\mu}{\mu+\lambda_{m}}\Big{)}||\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}||^{2}$
where coefficients satisfy
$\widehat{C}_{1}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho}\Big{)}$ $=$ $\frac{1}{\mu}-\frac{LB}{\mu^{2}\rho}-\frac{LB^{2}}{2\rho^{2}}$
$\widehat{C}_{2}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{\mu}{\mu+\lambda_{m}}\Big{)}$ $=$ $\Big{(}\frac{L^{2}}{\mu^{2}\rho}+\frac{L+\mu}{\mu^{2}}+\frac{L(L+\rho)B}{\rho^{2}}\Big{)}\frac{\mu}{\mu+\lambda_{m}}$
$\widehat{C}_{3}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{\mu}{\mu+\lambda_{m}}\Big{)}$ $=$ $\frac{L(L+\rho)^{2}}{2\rho^{2}}\frac{\mu^{2}}{(\mu+\lambda_{m})^{2}}$
Lemma D.2.
Let $\mathcal{S}_{m}$ be the set of $K$ randomly selected local devices got updated at iterations $12m-$ 1 and $2m$, and $\mathbb{E}_{\mathcal{S}_{m}}[\cdot]$ be expectation with respect to the choice of devices. Then we have
$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]$$\leq$ $f(\widetilde{\omega}^{2m+1})+\widetilde{C}_{1}\Big{(}B,L,\frac{1}{K},\frac{1}{\rho}\Big{)}||\nabla f(\overline{\omega}^{2m-1})||^{2}$
$+\widetilde{C}_{2}\Big{(}B,L,\frac{1}{K},\frac{1}{\rho},\frac{\mu}{\mu+\lambda_{m}}\Big{)}||\nabla f(\overline{\omega}^{2m-1})||\cdot||\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}||$
$+\widetilde{C}_{3}\Big{(}B,L,\frac{1}{K},\frac{1}{\rho},\frac{\mu}{\mu+\lambda_{m}}\Big{)}||\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}||^{2}$
where coefficients satisfy
$\widetilde{C}_{1}\Big{(}B,L,\frac{1}{K},\frac{1}{\rho}\Big{)}$ $=$ $\frac{2B^{2}}{K\rho^{2}}+\frac{2LB+\rho}{\rho}\sqrt{\frac{2}{K}}\frac{B}{\rho}$
$\widetilde{C}_{2}\Big{(}B,L,\frac{1}{K},\frac{1}{\rho},\frac{\mu}{\mu+\lambda_{m}}\Big{)}$ $=$ $\Big{(}\frac{4LB}{K\rho^{2}}+\frac{2LB+\rho}{\rho}\sqrt{\frac{2}{K}}\frac{L}{\rho}+2L\frac{L+\rho}{\rho}\sqrt{\frac{2}{K}}\frac{B}{\rho}\Big{)}\cdot\frac{\mu}{\mu+\lambda_{m}}$
$\widetilde{C}_{3}\Big{(}B,L,\frac{1}{K},\frac{1}{\rho},\frac{\mu}{\mu+\lambda_{m}}\Big{)}$ $=$ $\Big{(}\frac{2}{K}\frac{L^{2}}{\rho^{2}}+2L\frac{L+\rho}{\rho}\sqrt{\frac{2}{K}}\frac{L}{\rho}\Big{)}\cdot\Big{(}\frac{\mu}{\mu+\lambda_{m}}\Big{)}^{2}$
Appendix E Proofs
Proof of Lemma D.1
Proof.
Since local functions $F_{i}$ are $L$-Lipschitz smooth, at iteration $12m-$ 1, we have
$f(\widetilde{\omega}^{2m+1})$
$\leq f(\overline{\omega}^{2m-1})+\langle\nabla f(\overline{\omega}^{2m-1}),\widetilde{\omega}^{2m+1}-\overline{\omega}^{2m-1}\rangle+\frac{L}{2}||\widetilde{\omega}^{2m+1}-\overline{\omega}^{2m-1}||^{2}$
$=f(\overline{\omega}^{2m-1})+\langle\nabla f(\overline{\omega}^{2m-1}),-\frac{1}{\mu}\nabla f(\overline{\omega}^{2m-1})+\Phi^{2m+1}\rangle+\frac{L}{2}||\widetilde{\omega}^{2m+1}-\overline{\omega}^{2m-1}||^{2}$
$\leq f(\overline{\omega}^{2m-1})-\frac{1}{\mu}||\nabla f(\overline{\omega}^{2m-1})||^{2}+\frac{1}{\mu}||\nabla f(\overline{\omega}^{2m-1})||\cdot||\Phi^{2m+1}||+\frac{L}{2}||\widetilde{\omega}^{2m+1}-\overline{\omega}^{2m-1}||^{2}$
where
$\Phi^{2m+1}$ $=$ $\frac{1}{\mu}\nabla f(\overline{\omega}^{2m-1})+\widetilde{\omega}^{2m+1}-\overline{\omega}^{2m-1}=\mathbb{E}_{i}\Big{[}\frac{1}{\mu}\nabla F_{i}(\overline{\omega}^{2m-1})+\omega_{i}^{2m+1}-\overline{\omega}^{2m-1}\Big{]}$ (6)
By first-order condition, the following holds at $(2m+1)$-th iteration:
$\omega_{i}^{2m+1}-\overline{\omega}^{2m-1}=-\frac{1}{\mu}\nabla F_{i}(\omega_{i}^{2m+1})+\overline{\omega}^{2m}-\overline{\omega}^{2m-1}$
Plug into Eqn. (6), we have
$\Phi^{2m+1}$ $=$ $\mathbb{E}_{i}\Big{[}\frac{1}{\mu}\Big{(}\nabla F_{i}(\overline{\omega}^{2m-1})-\nabla F_{i}(\omega_{i}^{2m+1})\Big{)}+\overline{\omega}^{2m}-\overline{\omega}^{2m-1}\Big{]}$
By $L$-Lipschitz smoothness and Jensen’s inequality, we have
$||\Phi^{2m+1}||$$\leq$ $\mathbb{E}_{i}\Big{[}\frac{1}{\mu}||\nabla F_{i}(\overline{\omega}^{2m-1})-\nabla F_{i}(\omega_{i}^{2m+1})||\Big{]}+||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||$
$\leq$ $\mathbb{E}_{i}\Big{[}\frac{L}{\mu}||\overline{\omega}^{2m-1}-\omega_{i}^{2m+1}||\Big{]}+||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||$
$\leq$ $\mathbb{E}_{i}\Big{[}\frac{L}{\mu}||\omega_{i}^{2m+1}-\overline{\omega}^{2m}||\Big{]}+\frac{L+\mu}{\mu}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||$
Since $h_{i}$ are $\rho$-strongly convex, $F_{i}$ is $L$-Lipschitz smooth, and $\omega_{i}^{2m+1}=\arg\min_{\omega}h_{i}(\omega;\overline{\omega}^{2m})$ we have
$||\omega_{i}^{2m+1}-\overline{\omega}^{2m}||$$\leq$ $\frac{1}{\rho}||\nabla h_{i}(\omega_{i}^{2m+1};\overline{\omega}^{2m})-\nabla h_{i}(\overline{\omega}^{2m};\overline{\omega}^{2m})||=\frac{1}{\rho}||0-\nabla F_{i}(\overline{\omega}^{2m})||$ (8)
$\leq$ $\frac{L}{\rho}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||+\frac{1}{\rho}||\nabla F_{i}(\overline{\omega}^{2m-1})||$
Plug in Eqn. (E),
$||\Phi^{2m+1}||$$\leq$ $\frac{L}{\mu\rho}\mathbb{E}_{i}[||\nabla F_{i}(\overline{\omega}^{2m-1})||]+\Big{(}\frac{L^{2}}{\mu\rho}+\frac{L+\mu}{\mu}\Big{)}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||$
Consider the following term
$||\widetilde{\omega}^{2m+1}-\overline{\omega}^{2m-1}||$ $=$ $||\mathbb{E}_{i}[\omega_{i}^{2m+1}]-\overline{\omega}^{2m-1}||\leq\mathbb{E}_{i}[||\omega_{i}^{2m+1}-\overline{\omega}^{2m-1}||]$
$\leq$ $\mathbb{E}_{i}[||\omega_{i}^{2m+1}-\overline{\omega}^{2m}||+||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||]$
$\leq$ $\frac{L+\rho}{\rho}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||+\frac{1}{\rho}\mathbb{E}_{i}\Big{[}||\nabla F_{i}(\overline{\omega}^{2m-1})||\Big{]}$
Because $F_{i}$ is $B$-dissimilar, we have
$\mathbb{E}_{i}\Big{[}||\nabla F_{i}(\overline{\omega}^{2m-1})||\Big{]}\leq B||\nabla f(\overline{\omega}^{2m-1})||$
Therefore,
$f(\widetilde{\omega}^{2m+1})$
$\leq f(\overline{\omega}^{2m-1})-\frac{1}{\mu}||\nabla f(\overline{\omega}^{2m-1})||^{2}+\frac{1}{\mu}||\nabla f(\overline{\omega}^{2m-1})||\cdot||\Phi^{2m+1}||+\frac{L}{2}||\widetilde{\omega}^{2m+1}-\overline{\omega}^{2m-1}||^{2}$
$\leq f(\overline{\omega}^{2m-1})-\Big{(}\frac{1}{\mu}-\frac{LB}{\mu^{2}\rho}-\frac{LB^{2}}{2\rho^{2}}\Big{)}||\nabla f(\overline{\omega}^{2m-1})||^{2}$
$+\Big{(}\frac{L^{2}}{\mu^{2}\rho}+\frac{L+\mu}{\mu^{2}}+\frac{L(L+\rho)B}{\rho^{2}}\Big{)}||\nabla f(\overline{\omega}^{2m-1})||\cdot||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||$
$+\frac{L(L+\rho)^{2}}{2\rho^{2}}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||^{2}$
After applying first-order approximation in even iterations, we have
$\overline{\omega}^{2m}-\overline{\omega}^{2m-1}=\frac{\mu}{\mu+\lambda_{m}}(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2})$
Therefore,
$f(\widetilde{\omega}^{2m+1})$$\leq$ $f(\overline{\omega}^{2m-1})-\Big{(}\frac{1}{\mu}-\frac{LB}{\mu^{2}\rho}-\frac{LB^{2}}{2\rho^{2}}\Big{)}||\nabla f(\overline{\omega}^{2m-1})||^{2}$
$+\Big{(}\frac{L^{2}}{\mu^{2}\rho}+\frac{L+\mu}{\mu^{2}}+\frac{L(L+\rho)B}{\rho^{2}}\Big{)}\frac{\mu}{\mu+\lambda_{m}}||\nabla f(\overline{\omega}^{2m-1})||\cdot||\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}||$
$+\frac{L(L+\rho)^{2}}{2\rho^{2}}\frac{\mu^{2}}{(\mu+\lambda_{m})^{2}}||\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}||^{2}$
The Lemma D.1 is proved. ∎
Proof of Lemma D.2
Proof.
Because local function $F_{i}$ is $L$-Lipschitz smooth, $f$ is local Lipschitz continuous.
$f(\overline{\omega}^{2m+1})\leq f(\widetilde{\omega}^{2m+1})+L_{0}||\overline{\omega}^{2m+1}-\widetilde{\omega}^{2m+1}||$
where $L_{0}$ is the local Lipschitz continuouty constant. Moreover, we have
$L_{0}\leq||\nabla f(\overline{\omega}^{2m-1})||+L(||\widetilde{\omega}^{2m+1}-\overline{\omega}^{2m-1}||+||\overline{\omega}^{2m+1}-\overline{\omega}^{2m-1}||)$
Therefore,
$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]$
$\leq f(\widetilde{\omega}^{2m+1})$
$+\mathbb{E}_{\mathcal{S}_{m}}\Big{[}\Big{(}||\nabla f(\overline{\omega}^{2m-1})||+L(||\widetilde{\omega}^{2m+1}-\overline{\omega}^{2m-1}||+||\overline{\omega}^{2m+1}-\overline{\omega}^{2m-1}||)\Big{)}||\overline{\omega}^{2m+1}-\widetilde{\omega}^{2m+1}||\Big{]}$
$=f(\widetilde{\omega}^{2m+1})+\Big{(}||\nabla f(\overline{\omega}^{2m-1})||+L||\widetilde{\omega}^{2m+1}-\overline{\omega}^{2m-1}||\Big{)}\cdot\mathbb{E}_{\mathcal{S}_{m}}[||\overline{\omega}^{2m+1}-\widetilde{\omega}^{2m+1}||]$
$+L\mathbb{E}_{\mathcal{S}_{m}}\Big{[}||\overline{\omega}^{2m+1}-\overline{\omega}^{2m-1}||\cdot||\overline{\omega}^{2m+1}-\widetilde{\omega}^{2m+1}||\Big{]}$
$\leq f(\widetilde{\omega}^{2m+1})+\Big{(}||\nabla f(\overline{\omega}^{2m-1})||+2L||\widetilde{\omega}^{2m+1}-\overline{\omega}^{2m-1}||\Big{)}\cdot\mathbb{E}_{\mathcal{S}_{m}}[||\overline{\omega}^{2m+1}-\widetilde{\omega}^{2m+1}||]$
$+L\mathbb{E}_{\mathcal{S}_{m}}\Big{[}||\overline{\omega}^{2m+1}-\widetilde{\omega}^{2m+1}||^{2}\Big{]}$
When $K$ devices are randomly selected, by Eqn. (8), we have
$\mathbb{E}_{\mathcal{S}_{m}}\Big{[}||\overline{\omega}^{2m+1}-\widetilde{\omega}^{2m+1}||^{2}\Big{]}$
$\leq\frac{1}{K}\mathbb{E}_{i}\Big{[}||\omega_{i}^{2m+1}-\widetilde{\omega}^{2m+1}||^{2}\Big{]}\leq\frac{2}{K}\mathbb{E}_{i}\Big{[}||\omega_{i}^{2m+1}-\overline{\omega}^{2m}||^{2}\Big{]}$
$\leq\frac{2}{K}\mathbb{E}_{i}\Big{[}\frac{L^{2}}{\rho^{2}}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||^{2}+\frac{1}{\rho^{2}}||\nabla F_{i}(\overline{\omega}^{2m-1})||^{2}+\frac{2L}{\rho^{2}}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||\cdot||\nabla F_{i}(\overline{\omega}^{2m-1})||\Big{]}$
$\leq\frac{2}{K}\frac{L^{2}}{\rho^{2}}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||^{2}+\frac{2B^{2}}{K\rho^{2}}||\nabla f(\overline{\omega}^{2m-1})||^{2}+\frac{4LB}{K\rho^{2}}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||\cdot||\nabla f(\overline{\omega}^{2m-1})||$
$=\frac{2}{K}\Big{(}\frac{L}{\rho}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||+\frac{B}{\rho}||\nabla f(\overline{\omega}^{2m-1})||\Big{)}^{2}$
By Jensen’s inequality,
$\mathbb{E}_{\mathcal{S}_{m}}\Big{[}||\overline{\omega}^{2m+1}-\widetilde{\omega}^{2m+1}||\Big{]}$$\leq$ $\sqrt{\mathbb{E}_{\mathcal{S}_{m}}\Big{[}||\overline{\omega}^{2m+1}-\widetilde{\omega}^{2m+1}||^{2}\Big{]}}$
$=$ $\sqrt{\frac{2}{K}}\Big{(}\frac{L}{\rho}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||+\frac{B}{\rho}||\nabla f(\overline{\omega}^{2m-1})||\Big{)}$
By Eqn. (E),
$||\nabla f(\overline{\omega}^{2m-1})||+2L||\widetilde{\omega}^{2m+1}-\overline{\omega}^{2m-1}||\leq 2L\frac{L+\rho}{\rho}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||+\frac{2LB+\rho}{\rho}||\nabla f(\overline{\omega}^{2m-1})||$
Re-organize, we have
$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]$
$\leq f(\widetilde{\omega}^{2m+1})+\frac{2}{K}\frac{L^{2}}{\rho^{2}}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||^{2}+\frac{2B^{2}}{K\rho^{2}}||\nabla f(\overline{\omega}^{2m-1})||^{2}$
$+\frac{4LB}{K\rho^{2}}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||\cdot||\nabla f(\overline{\omega}^{2m-1})||$
$+\Big{(}2L\frac{L+\rho}{\rho}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||+\frac{2LB+\rho}{\rho}||\nabla f(\overline{\omega}^{2m-1})||\Big{)}\sqrt{\frac{2}{K}}\frac{L}{\rho}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||$
$+\Big{(}2L\frac{L+\rho}{\rho}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||+\frac{2LB+\rho}{\rho}||\nabla f(\overline{\omega}^{2m-1})||\Big{)}\sqrt{\frac{2}{K}}\frac{B}{\rho}||\nabla f(\overline{\omega}^{2m-1})||$
$=f(\widetilde{\omega}^{2m+1})+\Big{(}\frac{2}{K}\frac{L^{2}}{\rho^{2}}+2L\frac{L+\rho}{\rho}\sqrt{\frac{2}{K}}\frac{L}{\rho}\Big{)}\cdot\Big{(}\frac{\mu}{\mu+\lambda_{m}}\Big{)}^{2}||\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}||^{2}$
$+\Big{(}\frac{4LB}{K\rho^{2}}+\frac{2LB+\rho}{\rho}\sqrt{\frac{2}{K}}\frac{L}{\rho}+2L\frac{L+\rho}{\rho}\sqrt{\frac{2}{K}}\frac{B}{\rho}\Big{)}\cdot\frac{\mu}{\mu+\lambda_{m}}||\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}||\cdot||\nabla f(\overline{\omega}^{2m-1})||$
$+\Big{(}\frac{2B^{2}}{K\rho^{2}}+\frac{2LB+\rho}{\rho}\sqrt{\frac{2}{K}}\frac{B}{\rho}\Big{)}||\nabla f(\overline{\omega}^{2m-1})||^{2}$
Lemma D.2 is proved. ∎
Proof of Lemma 4.5
Proof.
Lemma 4.5 can be proved by combing Lemmas D.1 and D.2, where
$\boldsymbol{C_{1}}$ ::=:= $C_{1}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{1}{K}\Big{)}=\widehat{C}_{1}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho}\Big{)}-\widetilde{C}_{1}\Big{(}L,B,\frac{1}{K},\frac{1}{\rho}\Big{)}$ (10)
$=$ $\frac{1}{\mu}-\frac{LB}{\mu^{2}\rho}-\frac{LB^{2}}{2\rho^{2}}-\frac{2B^{2}}{K\rho^{2}}-\frac{2LB+\rho}{\rho}\sqrt{\frac{2}{K}}\frac{B}{\rho}$
$\boldsymbol{C_{2}}$ ::=:= $C_{2}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{1}{K},\frac{\mu}{\mu+\lambda_{m}}\Big{)}=\widehat{C}_{2}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{\mu}{\mu+\lambda_{m}}\Big{)}+\widetilde{C}_{2}\Big{(}L,B,\frac{1}{K},\frac{1}{\rho},\frac{\mu}{\mu+\lambda_{m}}\Big{)}$ (11)
$=$ $\Big{(}\frac{L^{2}}{\mu^{2}\rho}+\frac{L+\mu}{\mu^{2}}+\frac{L(L+\rho)B}{\rho^{2}}+\frac{4LB}{K\rho^{2}}+\frac{(4L^{2}B+\rho L(1+2B)}{\rho^{2}}\sqrt{\frac{2}{K}}\Big{)}\cdot\frac{\mu}{\mu+\lambda_{m}}$
$\boldsymbol{C_{3}}$ ::=:= $C_{3}\Big{(}L,\frac{1}{\mu},\frac{1}{\rho},\frac{1}{K},\frac{\mu}{\mu+\lambda_{m}}\Big{)}=\widehat{C}_{3}\Big{(}L,\frac{1}{\mu},\frac{1}{\rho},\frac{\mu}{\mu+\lambda_{m}}\Big{)}+\widetilde{C}_{3}\Big{(}L,\frac{1}{K},\frac{1}{\rho},\frac{\mu}{\mu+\lambda_{m}}\Big{)}$ (12)
$=$ $\Big{(}\frac{L(L+\rho)^{2}}{2\rho^{2}}+\frac{2}{K}\frac{L^{2}}{\rho^{2}}+2L\frac{L+\rho}{\rho}\sqrt{\frac{2}{K}}\frac{L}{\rho}\Big{)}\cdot\Big{(}\frac{\mu}{\mu+\lambda_{m}}\Big{)}^{2}$
Appendix F Proof of Theorem 4.6
This can be proved directed using Lemma 4.5 by averaging over all $M$ odd iterations.
Appendix G Proof of Theorem C.1
WLOG, consider the case when local device got updated in every iteration and the algorithm runs over $2M$ iterations in total. We will use the uppercase letters $X$ and lowercase letters $x$ to denote random variables and the corresponding realizations, and use $P_{X}(\cdot)$ to denote its probability distribution. To simplify the notations, we will drop the index $i$ as we are only concerned with one agent.
According to [27], for a mechanism $\mathscr{M}$ outputs $o$, with inputs $d$ and $\hat{d}$, let a random variable $c(o;\mathscr{M},d,\hat{d})=\log\frac{\text{Pr}(\mathscr{M}(d)=o)}{\text{Pr}(\mathscr{M}(\hat{d})=o)}$ denote the privacy loss at $o$, and
$\alpha_{\mathscr{M}}(\lambda)=\underset{d,\hat{d}}{\max}\log\mathbb{E}_{o\sim\mathscr{M}(d)}\{\exp(\lambda c(o;\mathscr{M},d,\hat{d}))\}$
For two neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ of agent $i$, by Lemma 5.1, the total privacy loss is only contributed by odd iterations. Thus, for any sequence of private (clipped) models $\widehat{\omega}^{t}$ generated by mechanisms $\{\mathscr{M}^{m}\}_{m=1}^{M}$ over $2M$ iterations, there is:
$c(\widehat{\omega}^{0:2M};\{\mathscr{M}^{m}\}_{m=1}^{M},\mathcal{D},\mathcal{D}^{\prime})$ $=$ $\log\frac{P_{\widehat{\Omega}^{0:2M}}(\widehat{\omega}^{0:2M}|\mathcal{D})}{P_{\widehat{\Omega}^{0:2M}}(\widehat{\omega}^{0:2M}|\mathcal{D}^{\prime})}$
$=$ $\sum_{m=0}^{M}\log\frac{P_{\widehat{\Omega}^{2m+1}}(\widehat{\omega}^{2m+1}|\mathcal{D},\widehat{\omega}^{0:2m})}{P_{\widehat{\Omega}^{2m+1}}(\widehat{\omega}^{2m+1}|\mathcal{D}^{\prime},\widehat{\omega}^{0:2m})}+\log\frac{P_{\widehat{\Omega}^{0}}(\widehat{\omega}^{0}|\mathcal{D})}{P_{\widehat{\Omega}^{0}}(\widehat{\omega}^{0}|\mathcal{D}^{\prime})}$
$=$ $\sum_{m=0}^{M}c(\widehat{\omega}^{2m+1};\mathscr{M}^{m},\widehat{\omega}^{0:2m},\mathcal{D},\mathcal{D}^{\prime})$
where $\widehat{\omega}^{0:t}=\{\widehat{\omega}^{\tau}\}_{\tau=0}^{t}$ and $\widehat{\Omega}^{t}$ is random variable whose realization is $\widehat{\omega}^{t}$. Since $\widehat{\omega}^{0}$ is randomly generated, which is independent of dataset, we have $P_{\widehat{\Omega}^{0}}(\widehat{\omega}^{0}|\mathcal{D})=P_{\widehat{\Omega}^{0}}(\widehat{\omega}^{0}|\mathcal{D}^{\prime})$. Moreover, the following holds for any $\lambda$:
$\log\mathbb{E}_{\widehat{\omega}^{0:2M}}\{\exp(\lambda c(\widehat{\omega}^{0:2M};\{\mathscr{M}^{m}\}_{m=1}^{M},\mathcal{D},\mathcal{D}^{\prime}))\}$ (13)
$=$ $\log\mathbb{E}_{\widehat{\omega}^{0:2M}}\{\exp(\lambda\sum_{m=0}^{M}c(\widehat{\omega}^{2m+1};\mathscr{M}^{m},\widehat{\omega}^{0:2m},\mathcal{D},\mathcal{D}^{\prime})\}$
$=$ $\sum_{m=0}^{M}\log\mathbb{E}_{\widehat{\omega}^{2m+1}}\{\exp(\lambda c(\widehat{\omega}^{2m+1};\mathscr{M}^{m},\widehat{\omega}^{0:2m},\mathcal{D},\mathcal{D}^{\prime})\}$
Therefore, $\alpha_{\{\mathscr{M}^{m}\}_{m=1}^{M}}(\lambda)\leq\sum_{m=1}^{M}\alpha_{\mathscr{M}^{m}}(\lambda)$ also holds. First bound each individual $\alpha_{\mathscr{M}^{m}}(\lambda)$.
Consider two neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$. Private (clipped) model $\widehat{\omega}^{2m+1}$ is generated by mechanism $\mathscr{M}^{m}(D)=\xi({\omega}^{2m+1})+N=\frac{1}{|D|}\sum_{d\in D}\eta(d)+N$ with function $||\eta(\cdot)||_{2}\leq\tau$ and Gaussian noise $N\sim\mathcal{N}(0,\sigma^{2}\textbf{I})$. Without loss of generality, let $\mathcal{D}^{\prime}=\mathcal{D}\cup\{d_{n}\}$, $f(d_{n})=\pm\tau\textbf{e}_{1}$ and $\sum_{d\in\mathcal{D}}\eta(d)=\textbf{0}$. Then $\mathscr{M}^{m}(\mathcal{D})$ and $\mathscr{M}^{m}(\mathcal{D}^{\prime})$ are distributed identically except for the first coordinate and the problem can be reduced to one-dimensional problem.
$c(\widehat{\omega}^{2m+1};\mathscr{M}^{m},\widehat{\omega}^{0:2m},\mathcal{D},\mathcal{D}^{\prime})$ $=$ $\log\frac{P_{\widehat{\Omega}^{2m+1}}(\widehat{\omega}^{2m+1}|\mathcal{D},\widehat{\omega}^{0:2m})}{P_{\widehat{\Omega}^{2m+1}}(\widehat{\omega}^{2m+1}|\mathcal{D}^{\prime},\widehat{\omega}^{0:2m})}$
$=$ $\log\frac{P_{N}(n)}{P_{N}(n\pm\tau)}$
$\leq$ $\frac{\tau}{2|D|\sigma^{2}}(2|n|+\tau)~{}.$
where $n+\frac{1}{|D|}\sum_{d\in\mathcal{D}}\eta(d)=\widehat{\omega}^{2m+1}$. Therefore,
$\alpha_{\mathscr{M}^{m}}(\lambda)$ $=$ $\log\mathbb{E}_{N\sim\mathcal{N}(0,\sigma^{2})}\{\exp(\lambda\frac{\tau}{2|D|\sigma^{2}}(2N+\tau))\}$
[…]
$=$ $\frac{\tau^{2}\lambda(\lambda+1)}{2|D|^{2}\sigma^{2}}~{}.$
$\alpha_{\{\mathscr{M}^{m}\}_{m=1}^{M}}(\lambda)$$\leq$ $\sum_{m=1}^{M}\alpha_{\mathscr{M}^{m}}(\lambda)=\frac{M\tau^{2}\lambda(\lambda+1)}{2|D|^{2}\sigma^{2}}$
Use the tail bound [Theorem 2, [27]], for any $\varepsilon\geq\frac{M\tau^{2}}{2|D|^{2}\sigma^{2}}$, the algorithm is $(\varepsilon,\delta)$-differentially private for
$\delta$ $=$ $\underset{\lambda:\lambda\geq 0}{\min}~{}h(\lambda)=\underset{\lambda:\lambda\geq 0}{\min}\exp\Big{(}\frac{M\tau^{2}\lambda(\lambda+1)}{2|D|^{2}\sigma^{2}}-\lambda\varepsilon\Big{)}$
To find $\lambda^{*}=\underset{\lambda:\lambda\geq 0}{\text{argmin}}~{}h(\lambda)$, take derivative of $h(\lambda)$ and assign 0 gives the solution $\bar{\lambda}=\frac{\varepsilon|D|^{2}\sigma^{2}}{M\tau^{2}}-\frac{1}{2}\geq 0$, and $h^{\prime\prime}(\bar{\lambda})>0$, implies $\lambda^{*}=\bar{\lambda}$. Plug into (G) gives:
$\delta=\exp\left(\Big{(}\frac{M\tau^{2}}{4|D|^{2}\sigma^{2}}-\frac{\varepsilon}{2}\Big{)}\Big{(}\frac{\varepsilon|D|^{2}\sigma^{2}}{M\tau^{2}}-\frac{1}{2}\Big{)}\right)$ (14)
Similarly, for any $\delta\in[0,1]$, the algorithm is $(\varepsilon,\delta)$-differentially private for
$\varepsilon$ $=$ $\underset{\lambda:\lambda\geq 0}{\min}~{}h_{1}(\lambda)=\underset{\lambda:\lambda\geq 0}{\min}~{}\frac{M\tau^{2}(\lambda+1)}{2|D|^{2}\sigma^{2}}+\frac{1}{\lambda}\log\left(\frac{1}{\delta}\right)=2\sqrt{\frac{M\tau^{2}}{2|D|^{2}\sigma^{2}}\log(\frac{1}{\delta})}+\frac{M\tau^{2}}{2|D|^{2}\sigma^{2}}$
Proof of Theorem C.2
Proof.
WLOG, consider the case when local device got updated in every iteration and the algorithm runs over $2M$ iteration in total.
We will use the uppercase letters $X$ and lowercase letters $x$ to denote random variables and the corresponding realizations, and use $P_{X}(\cdot)$ to denote its probability distribution. To simplify the notations, we will drop the index $i$ as we are only concerned with one agent, and use $\omega^{t}$ to denote private output $\widehat{\omega}^{t}$.
For two neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ of agent $i$, by Lemma 5.1, the total privacy loss is only contributed by odd iterations. Thus, the ratio of joint probabilities (privacy loss) is given by:
$\frac{P_{\Omega^{0:2M}}(\omega^{0:2M}|\mathcal{D})}{P_{\Omega^{0:2M}}(\omega^{0:2M}|\mathcal{D}^{\prime})}=\frac{P_{\Omega^{0}}(\omega^{0}|\mathcal{D})}{P_{\Omega^{0}}(\omega^{0}|\mathcal{D}^{\prime})}\cdot\prod_{m=0}^{M}\frac{P_{\Omega^{2m+1}}(\omega^{2m+1}|\omega^{0:2m},\mathcal{D})}{P_{\Omega^{2m+1}}(\omega^{2m+1}|\omega^{0:2m},\mathcal{D}^{\prime})}$ (15)
where $\omega^{0:t}:=\{\omega^{s}\}_{s=1}^{t}$ and $\Omega^{t}$ denotes random variable of $\omega^{t}$. Since $\omega^{0}$ is randomly generated, which is independent of dataset. We have $P_{\Omega^{0}}(\omega^{0}|\mathcal{D})=P_{\Omega^{0}}(\omega^{0}|\mathcal{D}^{\prime})$.
Consider the $(2m+1)$-th iteration, by first-order condition, we have:
$n^{m}=-\nabla F_{i}(\omega^{2m+1};\mathcal{D})-\mu(\omega^{2m+1}-\overline{\omega}^{2m}):=g(\omega^{2m+1};\mathcal{D})$
Given $\omega^{0:2m}$, $n^{m}$ and $\omega^{2m+1}$ will be bijective and the relation is captured by a one-to-one mapping $g:\mathbb{R}^{d}\to\mathbb{R}^{d}$ defined above. By Jacobian transformation, we have
$P_{\Omega^{2m+1}}(\omega^{2m+1}|\omega^{0:2m},\mathcal{D})=P_{N^{m}}(g(\omega^{2m+1};\mathcal{D}))\cdot|\text{det}(\textbf{J}(g(\omega^{2m+1};\mathcal{D})))|$
Therefore,
$\frac{P_{\Omega^{2m+1}}(\omega^{2m+1}|\omega^{0:2m},\mathcal{D})}{P_{\Omega^{2m+1}}(\omega^{2m+1}|\omega^{0:2m},\mathcal{D}^{\prime})}=\frac{P_{N^{m}}(g(\omega^{2m+1};\mathcal{D}))}{P_{N^{m}}(g(\omega^{2m+1};\mathcal{D}^{\prime}))}\cdot\frac{|\text{det}(\textbf{J}(g(\omega^{2m+1};\mathcal{D})))|}{|\text{det}(\textbf{J}(g(\omega^{2m+1};\mathcal{D}^{\prime})))|}$
Let $n^{m}:=g(\omega^{2m+1};\mathcal{D})$, $n^{m^{\prime}}:=g(\omega^{2m+1};\mathcal{D}^{\prime})$ be noise vectors that result in output $\omega^{2m+1}$ under neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ respectively. WLOG, let $d_{1}\in\mathcal{D}$ and $d^{\prime}_{1}\in\mathcal{D}^{\prime}$ be the data pints in two datasets that are different, and $\mathcal{D}\setminus d_{1}=\mathcal{D}^{\prime}\setminus d^{\prime}_{1}$. Because noise vector $N^{m}\sim\exp(-\alpha^{m}||n^{m}||)$, we have,
$\frac{P_{N^{m}}(g(\omega^{2m+1};\mathcal{D}))}{P_{N^{m}}(g(\omega^{2m+1};\mathcal{D}^{\prime}))}$$\leq$ $\exp(\alpha^{m}||n^{m}-n^{m^{\prime}}||)=\exp(\alpha^{m}||\nabla F_{i}(\omega^{2m+1};\mathcal{D}^{\prime})-\nabla F_{i}(\omega^{2m+1};\mathcal{D})||)$ (16)
$=$ $\exp\Big{(}\frac{\alpha^{m}}{|\mathcal{D}|}||\nabla F_{i}(\omega^{2m+1};d^{\prime}_{1})-\nabla F_{i}(\omega^{2m+1};d_{1})||\Big{)}\leq\exp\Big{(}\frac{2\alpha^{m}u_{1}}{|\mathcal{D}|}\Big{)}$
Jacobian matrix
$\textbf{J}(g(\omega^{2m+1};\mathcal{D})))=-\nabla^{2}F_{i}(\omega^{2m+1};\mathcal{D})-\mu\textbf{I}_{d}:=A$ (17)
Further define matrix
$A_{\Delta}=\textbf{J}(g(\omega^{2m+1};\mathcal{D}^{\prime})))-A=\frac{1}{|\mathcal{D}|}\Big{(}\nabla^{2}F_{i}(\omega^{2m+1};d_{1})-\nabla^{2}F_{i}(\omega^{2m+1};d^{\prime}_{1})\Big{)}$
Then
$\frac{|\text{det}(\textbf{J}(g(\omega^{2m+1};\mathcal{D})))|}{|\text{det}(\textbf{J}(g(\omega^{2m+1};\mathcal{D}^{\prime})))|}=\frac{|\text{det}(A)|}{|\text{det}(A_{\Delta}+A)|}=\frac{1}{|\text{det}(I+A^{-1}A_{\Delta})|}=\frac{1}{|\prod_{k=1}^{r}(1+\lambda_{k}(A^{-1}A_{\Delta}))|}$
where $\lambda_{k}(A^{-1}A_{\Delta})$ denotes the $k$-th largest eigenvalue of matrix $A^{-1}A_{\Delta}$. Under generalized linear models, $A_{\Delta}$ has rank at most 2. Because $-\frac{u_{2}}{|\mathcal{D}|\mu}\leq\lambda_{k}(A^{-1}A_{\Delta})\leq\frac{u_{2}}{|\mathcal{D}|\mu}$ and $\mu,u_{2},|\mathcal{D}|$ satisfy $\frac{u_{2}}{|\mathcal{D}|\mu}\leq 0.5$, we have,
$\frac{|\text{det}(\textbf{J}(g(\omega^{2m+1};\mathcal{D})))|}{|\text{det}(\textbf{J}(g(\omega^{2m+1};\mathcal{D}^{\prime})))|}\leq\frac{1}{|1-\frac{u_{2}}{|\mathcal{D}|\mu}|^{2}}=\exp(-2\ln(1-\frac{u_{2}}{|\mathcal{D}|\mu}))\leq\exp\Big{(}\frac{2.8u_{2}}{|\mathcal{D}|\mu}\Big{)}$ (18)
where the last inequality holds because $-\ln(1-x)<1.4x$, $\forall x\in[0,0.5]$.
Combine Eqn. (15), (18) and (16), we have
$\frac{P_{\Omega^{0:2M}}(\omega^{0:2M}|\mathcal{D})}{P_{\Omega^{0:2M}}(\omega^{0:2M}|\mathcal{D}^{\prime})}$$\leq$ $\prod_{m=0}^{M}\exp\Big{(}\frac{2\alpha^{m}u_{1}}{|\mathcal{D}|}\Big{)}\cdot\exp\Big{(}\frac{2.8u_{2}}{|\mathcal{D}|\mu}\Big{)}$
$=$ $\exp\Big{(}\sum_{m=0}^{M}\frac{2\alpha^{m}u_{1}\mu+2.8u_{2}}{|\mathcal{D}|\mu}\Big{)}$
Theorem C.1 is proved. ∎
Appendix H Experiments
H.1 Details of Datasets
Table 2: Details of datasets. Numbers in parentheses represent the amount of test data. All of the numbers round to integer.
Dataset Samples # of device Samples per device
mean stdev
Synthetic iid 6726(683) 30 224 166
0,0 13791(1395) 30 460 841
0.5,0.5 8036(818) 30 268 410
1,1 10493(1063) 30 350 586
Femnist 16421(1924) 50 328 273
Sent140 32299(8484) 52 621 105
Synthetic. The synthetic data is generated using the same method in [2]. We briefly describe the generating steps here. For each device k, $y_{k}$ is computed from a softmax function $y_{k}=\text{argmax}(\text{softmax}(W_{k}x_{k}+b_{k}))$. $W_{k}$ and $b_{k}$ are drawn from the same Gaussian distribution with mean $u_{k}$ and variance $1$, where $u_{k}\in N(0;\beta)$. $x_{k}\in N(v_{k};\Sigma)$. $v_{k}$ is drawn from a Gaussian distribution with mean $B_{k}\in\mathcal{N}(0,\gamma)$ and variance $1$. $\Sigma$ is diagonal with $\sum j,j=j^{-1.2}$. In such a setting, $\beta$ controls how many local models differ from each other and $\gamma$ controls how much local data at each device differs from that of other devices.
In our experiment, we take $k=30$, $x\in\mathcal{R}^{20}$, $W\in\mathcal{R}^{10*20}$, $b\in\mathcal{R}^{10}$. We generate 4 datasets in total. They’re Synthetic(iid) Synthetic(0,0) with $\beta=0$ and $\gamma=0$, Synthetic(0.5,0.5) with $\beta=0.5$ and $\gamma=0.5$ and Synthetic(1,1) with $\beta=1$ and $\gamma=1$.
Femnist: Similar with [2], we subsample 10 lower case characters (‘a’-‘j’) from EMNIST [45] and distribute 5 classes to each device. There are 50 devices in total. The input is 28*28 image.
Sent140: A text sentiment analysis task on tweets [46]. The input is sequence of length 25 and output is the probabilities of 2 classes.
A brief summary of dataset can be found in Table 2.
H.2 Details of the Model
Classifier and Loss function. In our experiment, we use logistic regression for Synthetic, Multilayer perceptron and CNN for Femnist, Stacked LSTM for Sent140. Note that without privacy concern, any classifier and loss function can be plugged in to Upcycled-FL. However, if we adopt objective perturbation as privacy protection, the loss function should also satisfy assumptions in Theorem C.2. Thus log loss $\mathscr{L}(z)=log(1+\exp(-z))$ is used. It’s not hard to verify that this loss function satisfies assumptions in Theorem C.2 by taking first and second order derivatives($\left|\mathscr{L}^{\prime}\right|\leq u_{1}=1\text{ and }\mathscr{L}^{\prime\prime}\leq u_{2}=\frac{1}{4}$). For Multilayer perceptron, we use a two layer network with hidden dimension 14*14. For CNN, a 3*3 convolutional layer with 3 channels is adopted, following by two linear layers(followed by activation and dropout) with dimension 1000 and 100. For Stacked LTSM, we use 2 layers with 128 dimension Sent140. The word embedding dimension of Sent140 is 300.
Implementation and Environment. Our code is implemented in Pytorch 1.11 [49]. We employ SGD as local optimizer, with momentum 0.5, and set number of epochs E to 20 at each iteration $m$. Learning rate is tuned to 0.05. We run the model on a single GTX 1080 Ti.
H.3 CNN on Femnist
Figure 4: Loss and accuracy on Femnist using CNN. Trained with fraction=0.3, straggler=90%. $\mu=0.1,\lambda=0.04$. Training(testing) accuracy is Upcycled-FL 96.36%(96.21%), FedProx 95.63%(95.74%).
H.4 Convergence on all datasets
H.4.1 90% Straggler
Figure 5: Loss and accuracy on synthetic datasets. Trained with fraction=0.3, straggler=90%. (a) Synthetic(iid); (b) Synthetic(0,0); (c) Synthetic(0.5,0.5); (d) Synthetic(1,1). $\mu=0.1,\lambda=0.04$ for Synthetic(iid), Synthetic(0,0), $\mu=1,\lambda=0.42$ for Synthetic(0.5,0.5), Synthetic(1,1).
(a) $\mu=0.1$
(b) $\mu=0.3$
(c) $\mu=0.5$
Figure 6: Loss and accuracy on Femnist. Trained with fraction=0.3, straggler=90%. $\lambda=0.04,0.12,0.21$ is set based on the value of $\mu=0.1,0.3,0.5$, respectively.
(a) $\mu=0.1$
(b) $\mu=0.3$
Figure 7: Loss and accuracy on Sent140. Trained with fraction=0.3, straggler=90%. $\mu=0.1,\lambda=0.04$ for (a), $\mu=0.3,\lambda=0.12$ for (b).
H.4.2 30% Straggler
Figure 8: Loss and accuracy on synthetic datasets. Trained with fraction=0.3, straggler=30%. (a) Synthetic(iid); (b) Synthetic(0,0); (c) Synthetic(0.5,0.5); (d) Synthetic(1,1). $\mu=0.1,\lambda=0.04$ for Synthetic(iid), Synthetic(0,0), $\mu=1,\lambda=0.42$ for Synthetic(0.5,0.5), Synthetic(1,1).
H.5 Privacy on all datasets
H.5.1 Objective perturbation
  • Convergence
    Figure 9: Objective perturbation on synthetic datasets. Trained with fraction=1, straggler=0%. $\mu=0.1,\lambda=0.04$. (a) Synthetic(iid); (b) Synthetic(0,0); (c) Synthetic(0.5,0.5); (d) Synthetic(1,1).
  • Impact of $\alpha$
    Figure 10: Impact of $\alpha$ on Femnist. Trained with fraction=1, straggler=0%. $\mu=0.1,\ \lambda=0.04$.
H.5.2 Output perturbation
Figure 11: Output perturbation on synthetic datasets. Trained with fraction=1, straggler=0%. $\mu=0.1,\lambda=0.04$. (a) Synthetic(iid); (b) Synthetic(0,0); (c) Synthetic(0.5,0.5); (d) Synthetic(1,1).
Figure 12: Objective and output perturbation on Femnist and Sent140. (a) shows the result using objective perturbation on Femnist. The average accuracy of private Upcycled-FL is 71.04%(70.71%), of private FedProx is 19.72%(18.82%). (b) shows the result using output perturbation on Femnist. The average accuracy of private Upcycled-FL is 79.43%(79.63%), of private FedProx is 78.43%(78.36%). (c) shows the result using objective perturbation on Sent140. The average accuracy of private Upcycled-FL is 74.36%(72.38%), of private FedProx is 69.94%(68.02%). (d) shows the result using output perturbation on Sent140. The average accuracy of private Upcycled-FL is 74.42%(72.52%), of private FedProx is 62.10%(62.13%).All trained with fraction=1, straggler=0%, $\mu=0.1,\lambda=0.04$. For objective perturbation, $\delta=0.001$.

### v2

@@
-Federated Learning with Reduced Information Leakage and Computation
+algorithmic
 Abstract
@@
 As a result, it can be particularly difficult to balance the privacy-accuracy trade-off when designing privacy-preserving FL algorithms.
-In this paper, we introduce Upcycled-FL, a novel federated learning framework with first-order approximation applied at every even iteration.
-Under this framework, half of the FL updates incur no information leakage and require much less computation.
-We first conduct the theoretical analysis on the convergence (rate) of Upcycled-FL, and then apply perturbation mechanisms to preserve privacy.
-Experiments on real-world data show that Upcycled-FL consistently outperforms existing methods over heterogeneous data, and significantly improves privacy-accuracy trade-off while reducing 48% of the training time on average.
+This paper introduces Upcycled-FL, a simple yet effective strategy that applies first-order approximation at every even round of model update.
+Under this strategy, half of the FL updates incur no information leakage and require much less computational and transmission costs.
+We first conduct the theoretical analysis on the convergence (rate) of Upcycled-FL and then apply two perturbation mechanisms to preserve privacy.
+Extensive experiments on both synthetic and real-world data show that the Upcycled-FL strategy can be adapted to many existing FL frameworks and consistently improve the privacy-accuracy trade-off.*** Code available at https://github.com/osu-srml/Upcycled-FL.
 1 Introduction
-Federated learning has emerged as an important paradigm for learning models in a distributed fashion, whereby data is distributed across different clients and the goal is to jointly learn a model from the distributed data.
-This is facilitated by a central server and the model can be learned through iterative interactions between the central server and clients: at each iteration, each client performs certain computation using its local data; the local results are collected and aggregated by the central server; the aggregated result is then sent to clients to update local results; and so on till the learning task is deemed accomplished.
-Although each client’s data is not shared directly with the central server, their sensitive information may be inferred from the intermediate computations.
-Indeed, as the same data is repeatedly used during the iterative learning process, each client’s total information leakage accumulates substantially over time.
-For randomization-based privacy-preserving techniques such as differential privacy [1], the perturbation required for guaranteeing the privacy of each client also grows as the total information leakage increases.
-As a result, it can be particularly difficult to balance the privacy and accuracy trade-off in FL.
-This paper proposes a novel federated learning framework called Upcycled Federated Learning (Upcycled-FL)***The word “upcycle” refers to reusing material so as to create higher-quality things than the original., in which clients’ information leakage can be reduced such that it only occurs during half of the updates.
-This is attained by modifying the even iterations of a baseline FL algorithm with first-order approximation, which allows us to compute the update using existing results from previous iterations without using the client’s data.
-Moreover, the updates in even iterations only involve addition/subtraction operations on existing results at the central server; the computational costs and the training time can be reduced significantly.
-To preserve privacy, we can apply existing privacy mechanisms to Upcycled-FL.
-As an example, we consider differential privacy (DP) because it can (i) rigorously quantify the total privacy leakage for complex algorithms such as FL; (ii) defend against attackers regardless of their background knowledge; (iii) provide heterogeneous privacy guarantees for different clients.
-Since DP preserves the client’s privacy by randomizing the algorithm (e.g., adding noise to the output, objective, or gradient of local clients), Upcycled-FL by reducing the total information leakage, requires less randomization to attain a certain level of privacy.
-Therefore, the privacy-accuracy trade-off is also improved.
-We emphasize that the idea of upcycling is orthogonal to both the baseline FL algorithm, and the DP perturbation method.
-In this paper, we adopt FedProx [2] as the baseline FL algorithm and apply first-order approximation to it, but our approach can be applied to other FL frameworks, such as FedNova [3] and pFedMe [4].
+Federated learning (FL) has emerged as an important paradigm for learning models in a distributed fashion, whereby data is distributed across different clients and the goal is to jointly learn a model from the distributed data.
+This is facilitated by a central server and the model can be learned through iterative interactions between the central server and clients: at each iteration, each client performs certain computation using its local data; the updated local models are collected and aggregated by the server; the aggregated model is then sent back to clients for them to update local models; and so on till the learning task is deemed accomplished.
+Although each client’s data is not shared directly with the central server, there is a risk of information leakage that a third party may infer individual sensitive information from (intermediate) computational outcomes.
+Information leakage occurs whenever the client’s local gradients are shared with third parties.
+Importantly, the information leakage (or privacy loss) accumulates as data is repeatedly used during the iterative learning process: with more computational outcomes derived from individual data, third parties have more information to infer sensitive data and it poses higher privacy risks for individuals.
+An example is Huang et al.
+(2021), which shows that eavesdroppers can conduct gradient inversion attacks to recover clients’ data from the gradients.
+In this paper, we use differential privacy (DP) proposed by Dwork (2006), a de facto standard for preserving individual data privacy in data analysis, ranging from simple tasks such as data collection and statistical analysis (Zhang et al., 2022b; Ghosh and Roth, 2011; Khalili and Vakilinia, 2021; Amin et al., 2019; Liu et al., 2021; Khalili et al., 2021c, 2019) to complex machine learning and optimization tasks (Cai et al., 2024; Khalili et al., 2021b, a; Huang et al., 2020; Jagielski et al., 2019; Zhang et al., 2019b, 2018a, 2018b, c).
+Compared to other privacy preservation techniques, it can (i) rigorously quantify the total privacy leakage for complex algorithms such as FL; (ii) defend against attackers regardless of their background knowledge; (iii) provide heterogeneous privacy guarantees for different clients.
+To achieve a certain DP guarantee, we need to perturb FL algorithms (e.g., adding noise to the output, objective, or gradient of local clients) and the perturbation needed for achieving the privacy requirement of each client grows as the total information leakage increases.
+Because the added perturbation inevitably reduces algorithm accuracy, it can be difficult to balance the privacy and accuracy trade-off in FL.
+This paper proposes a novel strategy for FL called Upcycled Federated Learning (Upcycled-FL)The word “upcycle” refers to reusing material so as to create higher-quality things than the original., in which clients’ information leakage can be reduced such that it only occurs during half of the updates.
+This is attained by modifying the even iterations of a baseline FL algorithm with first-order approximation, which allows us to update the model using existing model parameters from previous iterations without using the client’s data.
+Moreover, the updates in even iterations only involve addition/subtraction operations on existing model parameters at the central server.
+Because Upcycled-FL doesn’t require local training and transmission in half of iterations, the transmission costs and the training time can be reduced significantly.
+It turns out that Upcycled-FL, by reducing the total information leakage, requires less perturbation to attain a certain level of privacy and can enhance the privacy-accuracy trade-off significantly.
+We emphasize that the idea of “upcycling information” is orthogonal to both the baseline FL algorithm and the DP perturbation method.
+It can be applied to any FL algorithms that involve local optimization at the clients.
+In this paper, we apply Upcycled-FL strategy to multiple existing FL algorithms and evaluate them on both synthetic and real-world datasets.
 For DP perturbation methods, we consider both output and objective perturbation as examples to quantify the privacy loss, while other DP methods can also be used.
-It is worth noting that although differentially private federated learning has been extensively studied in the literature, e.g., [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], all these algorithms need client’s local data to update the model and the information leakage inevitably occurs at every iteration.
-This is fundamentally different from this work, where we propose a novel method for FL that reduces information leakage and computation.
-In addition to private federated learning, many approaches were proposed in the DP literature to improve the privacy-accuracy trade-off (e.g., privacy amplification by sampling [21, 22, 23, 24, 25, 26, 27], leveraging non-private public data [28, 29], shuffling [30], using weaker privacy notion [31], using tighter privacy composition analysis tool [27]).
+It is worth noting that although differentially private federated learning has been extensively studied in the literature, e.g., (Asoodeh et al., 2021; Chuanxin et al., 2020; Zhang et al., 2022a; Zheng et al., 2021; Wang et al., 2020b; Kim et al., 2021; Zhang et al., 2019a; Baek et al., 2021; Wu et al., 2022; Girgis et al., 2021; Truex et al., 2020; Hu et al., 2020; Seif et al., 2020; Zhao et al., 2020; Wei et al., 2020; Triastcyn and Faltings, 2019), all these algorithms need client’s local data to update the model and the information leakage inevitably occurs at every iteration.
+This is fundamentally different from this work, where we propose a novel strategy that effectively reduces information leakage in FL.
+In addition to private federated learning, several approaches were proposed in the DP literature to improve the privacy-accuracy trade-off, e.g., privacy amplification by sampling (Balle et al., 2018; Beimel et al., 2014; Hu et al., 2021; Wang et al., 2019; Kasiviswanathan et al., 2011; Wang et al., 2015; Abadi et al., 2016), leveraging non-private public data (Avent et al., 2017; Papernot et al., 2016), shuffling (Úlfar Erlingsson et al., 2019), using weaker privacy notion (Bun and Steinke, 2016), using tighter privacy composition analysis tool (Abadi et al., 2016).
 However, none of these strategies affect the algorithmic properties of learning algorithms.
 By contrast, our method improves the privacy-accuracy trade-off by modifying the property of FL algorithm (i.e., reducing the total information leakage); this improvement on the algorithmic property is independent of the privacy notion/mechanism or the analysis method.
-More related work is discussed in Appendix B
 Our main contributions are summarized as follows.
-  • We propose Upcycled-FL (Algorithm 2), a novel federated learning framework with reduced information leakage and computation.
-  • We conduct convergence (rate) analysis (Section 4, Theorem 4.6) and identify a sufficient condition for the convergence of Upcycled-FL.
-  • We apply output perturbation and objective perturbation to Upcycled-FL and conduct privacy analysis (Section 5, Theorem C.1).
-  • We evaluate the performance of (private) Upcycled-FL on real-world data (Section 6).
-The results show that Upcycled-FL, by reducing information leakage, outperforms existing algorithms over heterogeneous data and improves the accuracy-privacy trade-off.
-2 Problem Formulation and Preliminaries
+  • We propose Upcycled-FL (Algorithm 1), a novel strategy with reduced information leakage and computation that is broadly applicable to many existing FL algorithms.
+  • As an example, we apply our strategy to FedProx (Li et al., 2020) and conduct convergence (rate) analysis (Section 5, Theorem 5.6) where we identify a sufficient condition for the convergence of Upcycled-FL.
+  • As an example, we apply two differential privacy mechanisms (i.e., output perturbation and objective perturbation) to Upcycled-FL and conduct privacy analysis (Section 6, Theorem 6.2).
+  • We evaluate the effectiveness of Upcycled-FL on both synthetic and real data (Section 7).
+Extensive experiments show that Upcycled-FL can be adapted to many existing federated algorithms to achieve better performance; it effectively improves the accuracy-privacy trade-off by reducing information leakage.
+2 Related Work
+Differential privacy in federated learning.
+Differential privacy has been widely used in federated learning to provide privacy guarantees (Asoodeh et al., 2021; Chuanxin et al., 2020; Zhang et al., 2022a; Zheng et al., 2021; Wang et al., 2020b; Kim et al., 2021; Zhang et al., 2019a; Baek et al., 2021; Wu et al., 2022; Girgis et al., 2021; Truex et al., 2020; Hu et al., 2020; Seif et al., 2020; Zhao et al., 2020; Wei et al., 2020; Triastcyn and Faltings, 2019).
+For example, Zhang et al.
+(2022a) uses the Gaussian mechanism for a federated learning problem and propose an incentive mechanism to encourage users to share their data and participate in the training process.
+Zheng et al.
+(2021) introduces $f$-differential privacy, a generalized version of Gaussian differential privacy, and propose a federated learning algorithm satisfying this new notion.
+Wang et al.
+(2020b) proposes a new mechanism called Random Response with Priori (RRP) to achieve local differential privacy and apply this mechanism to the text data by training a Latent Dirichlet Allocation (LDA) model using a federated learning algorithm.
+Triastcyn and Faltings (2019) adapts the Bayesian privacy accounting method to the federated setting and propose joint accounting method for estimating client-level and instance-level privacy simultaneously and securely.
+Wei et al.
+(2020) presents a private scheme that adds noise to parameters at the random selected devices before aggregating and provides a convergence bound.
+Kim et al.
+(2021) combines the Gaussian mechanism with gradient clipping in federated learning to improve the privacy-accuracy tradeoff.
+Asoodeh et al.
+(2021) considers a different setting where only the last update is publicly released and the central server and other devices are assumed to be trustworthy.
+However, all these algorithms need client’s local data to update the model and the information leakage inevitably occurs at every iteration.
+This is fundamentally different from Upcycled-FL, which reuses existing results to update half of iterations and significantly reduces information leakage and computation.
+Tackling heterogeneity in federated learning.
+It’s worth mentioning that, Upcycled-FL also empirically outperforms existing baseline algorithms under device and statistical heterogeneity.
+In real-world scenarios, local data are often non-identically distributed across different devices; different devices are also often equipped with different specifications and computation capabilities.
+Such heterogeneity often causes instability in the model performance and leads to divergence.
+Many approaches have been proposed to tackle this issue in FL.
+For example, FedAvg (McMahan et al., 2017) uses a random selection of devices at each iteration to reduce the negative impact of statistical heterogeneity; however, it may fail to converge when heterogeneity increases.
+Other methods include FedProx (Li et al., 2020), a generalization and re-parameterization of FedAvg that adds a proximal term to the objective function to penalize deviations in the local model from the previous aggregation, and FedNova (Wang et al., 2020a) that re-normalizes local updates before updating to eliminate objective inconsistency.
+It turns out that Upcycled-FL exhibits superior performance in the presence of heterogeneity because gradients encapsulate information on data heterogeneity, the reusing of which leads to a boost in performance.
+3 Problem Formulation
 Consider an FL system consisting of a central server and a set $\mathcal{I}$ of clients.
-Each client $i$ has its local data $\mathcal{D}_{i}$ and they can be non-i.i.d across the clients.
+Each client $i$ has its local dataset $\mathcal{D}_{i}$ and these datasets can be non-i.i.d across the clients.
 The goal of FL is to learn a model $\omega\in\mathbb{R}^{d}$ from data $\cup_{i\in\mathcal{I}}\mathcal{D}_{i}$ by solving the following optimization:
@@
 Let $\omega_{i}^{t}$ be client $i$’s local model parameter at time $t$.
-In FL, the model is learned through an iterative process: at each time step $t$, (1) local computations: each active client updates its local model $\omega_{i}^{t}$ using its local data $\mathcal{D}_{i}$; (2) local models broadcasts: local models (or gradients) are then uploaded to the central server; (3) model aggregation: the central server aggregates results received from clients to update the global model parameter $\overline{\omega}^{t}=\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t}$; (4) model updating: the aggregated model is sent back to clients and is used for updating local models at $t+1$.
-In this work, we consider a baseline FL framework FedProx and apply the proposed method to it, while our method can also be used on other FL frameworks such as FedNova [3] and pFedMe [4].
-The details of FedProx are given in Algorithm 1.
-Algorithm 1 FedProx [2]
-1:  Input: $\mu>0$, $\{\mathcal{D}_{i}\}_{i\in\mathcal{I}}$, $\overline{\omega}^{0}$
-2:  for $t=1$ to $T$ do
-3:     The central server sends the current global model parameter $\overline{\omega}^{t}$ to all the clients.
-4:     A subset of clients get active and each active client updates its local model by finding (approximate) minimizer of local loss function:
-$\omega_{i}^{t+1}=\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{t}||^{2}.$
-5:     Each client sends its local model to server.
-6:     The central server updates the global model by aggregating all local models:
-$\overline{\omega}^{t+1}=\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t+1}.$
-7:  end for
+In FL, the model is learned through an iterative process: at each time step $t$, 1) local computations: each active client updates its local model $\omega_{i}^{t}$ using its local data $\mathcal{D}_{i}$; 2) local models broadcasts: local models (or gradients) are then uploaded to the central server; 3) model aggregation: the central server aggregates results received from clients to update the global model parameter $\overline{\omega}^{t}=\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t}$; 4) model updating: the aggregated model is sent back to clients and is used for updating local models at $t+1$.
 During the learning process, each client’s local computation is exposed to third parties at every iteration: its models/gradients need to be uploaded to the central server, and the global models calculated based on them are shared with all clients.
@@
 In this work, we consider differential privacy as the notion of privacy.
-Differential Privacy [1].
+Differential Privacy (Dwork, 2006).
 Differential privacy (DP) centers around the idea that the output of a certain computational procedure should be statistically similar given singular changes to the input, thereby preventing meaningful inference from observing the output.
 In FL, the information exposed by each client $i$ includes all intermediate computations $\{\omega_{i}^{t}\}_{t=1}^{T}$.
-Consider a randomized FL algorithm $\mathcal{A}(\cdot)$ that generates a sequence of private local models $\{\widehat{\omega}_{i}^{t}\}_{t=1}^{T}$, we say it satisfies DP for client $i$ over $T$ iterations if the following holds for any possible output $O\in\mathbb{R}^{d}\times\cdots\times\mathbb{R}^{d}$, and for any two neighboring local datasets $\mathcal{D}_{i}$, $\mathcal{D}^{\prime}_{i}$:
+Consider a randomized FL algorithm $\mathcal{A}(\cdot)$ that generates a sequence of private local models $\{\widehat{\omega}_{i}^{t}\}_{t=1}^{T}$, we say it satisfies $(\varepsilon,\delta)$-differential privacy for client $i$ over $T$ iterations if the following holds for any possible output $O\in\mathbb{R}^{d}\times\cdots\times\mathbb{R}^{d}$, and for any two neighboring local datasets $\mathcal{D}_{i}$, $\mathcal{D}^{\prime}_{i}$:
 $\Pr(\{\widehat{\omega}_{i}^{t}\}_{t=0}^{T}\in O|\mathcal{D}_{i})\leq\exp{(\varepsilon)}\cdot\Pr(\{\widehat{\omega}_{i}^{t}\}_{t=0}^{T}\in O|\mathcal{D}^{\prime}_{i})+\delta.$
@@
 Two datasets are neighboring datasets if they are different in at most one data point.
-3 Proposed Algorithm: Upcycled-FL
-Main idea: Fundamentally, the accumulation of information leakage over iterations stems from the fact that the client’s data $\mathcal{D}_{i}$ is used in every update.
-If the updates can be made without directly using this original data, but only from computational results that already exist, then the information leakage originating from these updates will be zero, and meanwhile the computational cost may be reduced significantly.
-Based on this idea, we propose Upcycled-FL, which modifies FedProx such that the earlier computations are repeatedly used for new updates.
-Upcycling Information: Next, we present Upcycled-FL and illustrate how the client’s total information leakage is reduced under this method.
-We modify FedProx by applying first-order approximation to even iterations.
-Specifically, at even iteration $2m$, we expand $F_{i}(\omega;\mathcal{D}_{i})$ at $\omega^{2m-1}_{i}$ (local model in the previous iteration).
-Based on Taylor series expansion, we have:
-$F_{i}(\omega;\mathcal{D}_{i})$ $=$ $F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})$
-$+\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}(\omega-\omega^{2m-1}_{i})$
-$+\mathcal{O}(||\omega-\omega^{2m-1}_{i}||^{2}).$
-We approximate term $\mathcal{O}(||\omega-\omega^{2m-1}_{i}||^{2})$ by $\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}$ for some constant $\lambda_{m}\geq 0$, which can be different for different iteration $2m$, then the even update of FedProx under such an approximation becomes:
-$\omega_{i}^{2m}$ $=$ $\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2}$ (2)
-$\approx$ $\arg\min_{\omega}F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})+\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}\omega$
-$+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2}$
-$=$ $\arg\min_{\omega}\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}\omega$
-$+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2}.$
-Note that in the above even update, the only term that depends on dataset $\mathcal{D}_{i}$ is $\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})$, which can be derived directly from the previous odd iteration.
+4 Proposed Method: Upcycled-FL
+Main idea.
+Fundamentally, the accumulation of information leakage over iterations stems from the fact that the client’s data $\mathcal{D}_{i}$ is used in every update.
+If the updates can be made without directly using this original data, but only from computational results that already exist, then the information leakage originating from these updates will be zero, and meanwhile, the computational cost may be reduced significantly.
+Based on this idea, we propose Upcycled-FL, which considers reusing the earlier computations in a new update and significantly reduces total information leakage and computational cost.
+Note that Upcycled-FL is not a specific algorithm but an effective strategy that can be used for any existing FL algorithms.
+Upcycling model update.
+Next, we present Upcycled-FL and illustrate how the client’s total information leakage is reduced under this method.
+For an FL system with the objective shown in Eqn.
+(1), the client $i$’s local objective function is given by $F_{i}(\omega;\mathcal{D}_{i})$.
+Under Upcycled-FL, we apply first-order approximation to $F_{i}(\omega;\mathcal{D}_{i})$ at even iterations during federated training (while odd updates remain fixed).
+Specifically, at $2m$-th iteration, we expand $F_{i}(\omega;\mathcal{D}_{i})$ at $\omega^{2m-1}_{i}$ (local model in the previous iteration).
+Based on the Taylor series expansion, we have:
+$F_{i}(\omega;\mathcal{D}_{i})$ $=$ $F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})+\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}(\omega-\omega^{2m-1}_{i})+\mathcal{O}(||\omega-\omega^{2m-1}_{i}||^{2})$ (2)
+$\approx$ $F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})+\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}(\omega-\omega^{2m-1}_{i})+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}$
+for some constant $\lambda_{m}\geq 0$ which may differ for different iteration $2m$.
+Then for an FL algorithm, its model update at $2m$-th iteration under Upcycled-FL strategy can be attained by replacing $F_{i}(\omega;\mathcal{D}_{i})$ with its approximation in Eqn.
+(2) while the updates at odd iterations remain the same.
+We illustrate this using the following two examples.
+Example 4.1 (FedAvg (McMahan et al., 2017) under Upcycled-FL strategy).
+In FedAvg, client $i$ at each iteration updates the local model by optimizing its local objective function, i.e., $\omega_{i}^{t}=\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i}),\forall t$.
+Under Upcycled-FL strategy, the client $i$’s updates become:
+$\omega_{i}^{t}=\begin{cases}\arg\min_{\omega}\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}\omega+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2},~{}~{}\text{ if }~{}~{}t=2m\\ \arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i}),~{}~{}\text{ if }~{}~{}t=2m-1\par\end{cases}$
+Example 4.2 (FedProx (Li et al., 2020) under Upcycled-FL strategy).
+In FedProx, a proximal term is added to the local objective function to stabilize the algorithm under heterogeneous clients (Algorithm 2), i.e., client $i$ at each iteration updates the local model $\omega_{i}^{t}=\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{t-1}||^{2},\forall t$.
+Under Upcycled-FL strategy, the client $i$’s updates become:
+$\omega_{i}^{t}=\begin{cases}\arg\min_{\omega}\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}\omega+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2},~{}~{}\text{ if }~{}~{}t=2m\\ \arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-2}||^{2},~{}~{}\text{ if }~{}~{}t=2m-1\end{cases}$ (3)
+Next, we demonstrate how the information is upcycled under the above idea. As an illustrating example, we focus on FedProx given in Example 4.2.
+Note that in the even update of Eqn.
+(3), the only term that depends on dataset $\mathcal{D}_{i}$ is $\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})$, which can be derived directly from the previous odd iteration.
 Specifically, according to the first-order condition, the following holds at odd iterations:
-$\omega_{i}^{2m-1}=\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-2}||^{2}$
 […]
-Plug $\nabla F_{i}(\omega_{i}^{2m-1};\mathcal{D}_{i})$ into even updates (2), we have
-$\omega_{i}^{2m}$$\approx$ $\arg\min_{\omega}\mu(\overline{\omega}^{2m-2}-\omega_{i}^{2m-1})^{T}\omega$ (4)
-$+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2}.$
-By first-order condition, even updates (4) can be reduced to:
+Plug $\nabla F_{i}(\omega_{i}^{2m-1};\mathcal{D}_{i})$ into even update of (3), we have the estimated update from the odd update:
+$\omega_{i}^{2m}$$\approx$ $\arg\min_{\omega}\mu(\overline{\omega}^{2m-2}-\omega_{i}^{2m-1})^{T}\omega+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2}.$ (5)
+where the approximately equals sign $\approx$” is due to the approximation in Eqn. (2). By first-order condition, even update (5) can be reduced to:
 $\omega_{i}^{2m}\approx\omega^{2m-1}_{i}+\frac{\mu}{\mu+\lambda_{m}}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right).$ (6)
 It turns out that with first-order approximation, dataset $\mathcal{D}_{i}$ is not used in the even updates.
-Furthermore, these updates can be conducted at the central server directly.
+Because these updates do not require access to the client’s data, these updates can be conducted at the central server directly.
 That is, the central server updates the global model by aggregating:
-$\overline{\omega}^{2m}$ $=$ $\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{2m}$
-$\approx$ $\sum_{i\in\mathcal{I}}p_{i}\omega^{2m-1}_{i}+\frac{\mu}{\mu+\lambda_{m}}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right)$
-$=$ $\overline{\omega}^{2m-1}+\frac{\mu}{\mu+\lambda_{m}}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right)$
-Therefore, even updates only involve addition/subtraction operations on the existing global models from previous iterations (i.e., $\overline{\omega}^{2m-1},\overline{\omega}^{2m-2}$): the computational cost and training time is reduced significantly.
-Note that the first-order approximation is only applied to even iterations, while the odd iterations should remain the same as FedProx to ensure Eqn.
+$\overline{\omega}^{2m}=\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{2m}\approx\sum_{i\in\mathcal{I}}p_{i}\omega^{2m-1}_{i}+\frac{\mu}{\mu+\lambda_{m}}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right)=\overline{\omega}^{2m-1}+\frac{\mu}{\mu+\lambda_{m}}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right)$
+Therefore, under Upcycled-FL strategy, even updates only involve addition/subtraction operations on the existing global models from previous iterations (i.e., $\overline{\omega}^{2m-1},\overline{\omega}^{2m-2}$) without the need for a local training epoch: both the computational cost and transmission cost are reduced significantly.
+Note that the first-order approximation is only applied to even iterations, while the odd iterations should remain as the origin to ensure Eqn.
 (4) holds.
@@
 In the following sections, we first analyze the convergence property of Upcycled-FL and then apply privacy mechanisms to satisfy differential privacy.
-Discussion: Indeed, if we view two consecutive steps (odd $12m-$ 1 followed by even $2m$) of Upcycled-FL as a single iteration $t$, then Upcycled-FL and FedProx will incur the same information leakage but will differ at the phase of global model aggregation.
-Specifically, instead of simply aggregating the global model by averaging over local updates (i.e., $\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t}$) as in FedProx, Upcycled-FL not only takes the average of local updates but also push the aggregation moving more toward the updating direction (i.e., $\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t}-\overline{\omega}^{t-1}$).
-In Figure 1, we illustrate the difference between FedProx and Upcycled-FL from the two perspectives: (i) Upcycled-FL can be regarded as reusing the intermediate updates of FedProx, compared to which the total information leakage is reduced; or (ii) Upcycled-FL can be regarded as FedProx with a different global aggregation rule, and the larger global update in Upcycled-FL accelerates the learning process with the same information leakage.
 (a) Upcycled-FL reuses the intermediate updates
 (b) Upcycled-FL uses different aggregation rule
-Figure 1: Consider Upcycled-FL from two perspectives
+Figure 1: Upcycled-FL can be considered from two perspectives: (i) it can be regarded as reusing the intermediate updates of local models to reduce the total information leakage; or (ii) it can be regarded as a global aggregation method with larger global update, which accelerates the learning process with the same information leakage under the same training iterations.
+Discussion.
+Indeed, if we view two consecutive steps (odd $12m-$ 1 followed by even $2m$) of Upcycled-FL as a single iteration $t$, then Upcycled-FedProx and FedProx will incur the same information leakage but will differ at the phase of global model aggregation, as shown in Figure 1(b).
+Specifically, instead of simply aggregating the global model by averaging over local updates (i.e., $\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t}$), Upcycled-FL not only takes the average of local updates but also pushes the aggregation moving more toward the updating direction (i.e., $\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t}-\overline{\omega}^{t-1}$).
+We present the difference between Upcycled-FL update strategy and regular aggregation strategy from these two perspectives in Figure 1.
+As Upcycled-FL only accesses client data at even iterations, it halves the communication cost compared to standard FL methods with the same number of iterations.
+Algorithm 1 Proposed aggregation strategy: Upcycled-FL
+1:  Input: $\lambda_{m}>0$, $\mu>0$, $\{\mathcal{D}_{i}\}_{i\in\mathcal{I}}$, $\overline{\omega}^{0}$
+2:  for $m=1$ to $M$ do
+3:    The central server sends the global model parameters $\overline{\omega}^{2m-2}$ to all the clients.
+4:    A subset of clients are selected to be active and each active client updates its local model by finding an (approximate) minimizer of the local objective function:
+$\omega_{i}^{2m-1}\leftarrow\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})~{}~{}~{}~{}\text{ or~{} { other local objective functions}}$
+5:    Clients send local models to the central server.
+6:    The central server updates the global model by aggregating all local models:
+$\overline{\omega}^{2m-1}$ $=$ $\textstyle\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{2m-1}$
+$\overline{\omega}^{2m}$ $=$ $\overline{\omega}^{2m-1}+\frac{\lambda_{m}}{2}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right)$
+7:  end for
 5 Convergence Analysis
-In this section, we analyze the convergence property of Upcycled-FL.
+In this section, we analyze the convergence of Upcycled-FL.
+For illustrating purposes, we focus on analyzing the convergence of Upcycled-FedProx.
 Note that we do not require local functions $F_{i}(\cdot)$ to be convex.
 Moreover, we consider practical settings where data are non-i.i.d across different clients.
-Similar to [2], we introduce a measure below to quantify the dissimilarity between clients in the federated network.
-Definition 4.1 ($B$-Dissimilarity [2]).
+Similar to Li et al.
+(2020), we introduce a measure below to quantify the dissimilarity between clients in the federated network.
+Definition 5.1 ($B$-Dissimilarity (Li et al., 2020)).
 The local loss function $F_{i}$ is $B$-dissimilar if $\forall\omega$, we have $\mathbb{E}[||\nabla F_{i}(\omega)||^{2}]\leq||\nabla f(\omega)||^{2}B^{2}$.
-where $\mathbb{E}[\cdot]$ denotes the expectation over clients (Eqn.
+where $\mathbb{E}[\cdot]$ denotes the expectation over clients (see Eqn.
 (1)).
@@
 $\forall i$, $h_{i}(\omega;\overline{\omega}^{t}):=F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{t}||^{2}$ are $\rho$-strongly convex.
-The above assumptions are fairly standard. They first appeared in [2] and are adopted in subsequent works such as [4, 32, 33].
+The above assumptions are fairly standard.
+They first appeared in Li et al.
+(2020) and are adopted in subsequent works such as T Dinh et al.
+(2020); Khaled et al.
+(2020); Pathak and Wainwright (2020).
 Note that strongly convex assumption is not on local objective $F_{i}(\omega;\mathcal{D}_{i})$, but the regularized function $F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{t}||^{2}$, i.e., the assumption can be satisfied by selecting a sufficiently large $\mu>0$.
-Indeed, as shown in Section 6, our algorithm still converges even when Assumption 4.2 and 4.4 do not hold (e.g., DNN).
-Next, we conduct the theoretical analysis on the convergence (rate) of Upcycled-FL.
-Algorithm 2 Proposed framework: Upcycled-FL
-1:  Input: $\lambda_{m}>0$, $\mu>0$, $\{\mathcal{D}_{i}\}_{i\in\mathcal{I}}$, $\overline{\omega}^{0}$
-2:  for $m=1$ to $M$ do
-3:     The central server sends the current global model parameter $\overline{\omega}^{2m-2}$ to all the clients.
-4:     A subset of clients are selected to be active and each active client updates its local model by finding (approximate) minimizer of local loss function:
-$\omega_{i}^{2m-1}\leftarrow\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-2}||^{2}.$
-5:     Clients send local models to the central server.
-6:     The central server updates the global model by aggregating all local models:
-$\overline{\omega}^{2m-1}$ $=$ $\textstyle\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{2m-1}$
-$\overline{\omega}^{2m}$ $=$ $\overline{\omega}^{2m-1}+\frac{\mu}{\mu+\lambda_{m}}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right)$
-7:  end for
+Indeed, as shown in Section 7, our algorithm still converges even when Assumption 5.2 and 5.4 do not hold (e.g., DNN).
+Next, we conduct the theoretical analysis on the convergence rate.
 Lemma 5.5.
-Let $\mathcal{S}_{m}$ be a set of $K$ randomly selected clients got updated (i.e., active clients) at iterations $12m-$ 1 and $2m$, and $\mathbb{E}_{\mathcal{S}_{m}}[\cdot]$ be the expectation with respect to the choice of clients.
-Then under Assumption 4.2 and 4.4, we have
-$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]$$\leq$ $f(\overline{\omega}^{2m-1})-\boldsymbol{C_{1}}||\nabla f(\overline{\omega}^{2m-1})||^{2}$
-$+\boldsymbol{C_{2}}\cdot h^{1}_{m}+\boldsymbol{C_{3}}\cdot h^{2}_{m},$
+Let $\mathcal{S}_{m}$ be a set of $K$ randomly selected clients which got updated (i.e., active clients) at iterations $12m-$ 1 and $2m$, and $\mathbb{E}_{\mathcal{S}_{m}}[\cdot]$ be the expectation with respect to the choice of clients.
+Then under Assumption 5.2 and 5.4, we have
+$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]$$\leq$ $f(\overline{\omega}^{2m-1})-\boldsymbol{C_{1}}||\nabla f(\overline{\omega}^{2m-1})||^{2}+\boldsymbol{C_{2}}\cdot h^{1}_{m}+\boldsymbol{C_{3}}\cdot h^{2}_{m},$
 where
@@
 $h^{2}_{m}$ ::=:= $||\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}||^{2}.$
-The details of the term $\boldsymbol{C_{1}}$, $\boldsymbol{C_{2}}$, $\boldsymbol{C_{3}}$ (expressed as functions of $L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{1}{K},\frac{\mu}{\lambda_{m}}$) are in Appendix E, Eqn.
+The details of term $\boldsymbol{C_{1}}$, $\boldsymbol{C_{2}}$, $\boldsymbol{C_{3}}$ (expressed as functions of $L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{1}{K},\frac{\mu}{\lambda_{m}}$) are in Appendix C, Eqn.
 (11)-(13).
@@
 It is easy to verify $\boldsymbol{C_{2}},\boldsymbol{C_{3}}\geq 0$.
-By rearranging and telescoping, we get the following convergence rate of Upcycled-FL.
-Theorem 4.6 (Convergence rate of Upcycled-FL).
-Under Assumption 4.2 and 4.4, if $\boldsymbol{C_{1}}>0$, we have
-$\min_{m\in[M]}\mathbb{E}[||\nabla f(\overline{\omega}^{2m-1})||^{2}]$
-$\leq$ $\frac{1}{M}\sum_{m=0}^{M}\mathbb{E}[||\nabla f(\overline{\omega}^{2m-1})||^{2}]$
+By rearranging and telescoping, we get the following convergence rate of Upcycled-FedProx.
+Theorem 5.6 (Convergence rate of Upcycled-FedProx).
+Under Assumption 5.2 and 5.4, if $\boldsymbol{C_{1}}>0$, we have
+$\min_{m\in[M]}\mathbb{E}\left[||\nabla f(\overline{\omega}^{2m-1})||^{2}\right]$$\leq$ $\frac{1}{M}\sum_{m=0}^{M}\mathbb{E}\left[||\nabla f(\overline{\omega}^{2m-1})||^{2}\right]$
 $\leq$ $\frac{f(\overline{\omega}^{0})-f(\overline{\omega}^{*})}{M\boldsymbol{C_{1}}}+\frac{\sum_{m=0}^{M}\boldsymbol{C_{2}}h^{1}_{m}}{M\boldsymbol{C_{1}}}+\frac{\sum_{m=0}^{M}\boldsymbol{C_{3}}h^{2}_{m}}{M\boldsymbol{C_{1}}},$
-where $\overline{\omega}^{0}$ and $\overline{\omega}^{*}$ denote the initial and the optimal global model parameters, and $\boldsymbol{C_{2}}$ and $\boldsymbol{C_{3}}$ are decreasing in $\frac{\lambda_{m}}{\mu}$.
-Theorem 4.6 implies that tunable $\mu$ and $\lambda_{m}$ are key hyper-parameters that control the convergence (rate) and robustness of Upcycled-FL.
+where $\overline{\omega}^{0}$ and $\overline{\omega}^{*}$ denote the initial and the optimal global model parameters, respectively.
+Both terms $\boldsymbol{C_{2}}$ and $\boldsymbol{C_{3}}$ are decreasing in $\frac{\lambda_{m}}{\mu}$.
+Theorem 5.6 implies that tunable $\mu$ and $\lambda_{m}$ are key hyper-parameters that control the convergence (rate) and robustness of Upcycled-FedProx.
 Recall that $\mu$ penalizes the deviation of local model $\omega_{i}^{2m}$ from global aggregated model $\overline{\omega}^{2m-1}$, while $\lambda_{m}$ penalizes the update of local model $\omega_{i}^{2m}$ from its previous update $\omega_{i}^{2m-1}$.
@@
 clients).
-Although $\boldsymbol{C_{1}}$ may become negative as $B$ increases, the experiments in Section 6 show that Upcycled-FL can still converge when data is highly heterogeneous.
+Although $\boldsymbol{C_{1}}$ may become negative as $B$ increases, the experiments in Section 7 show that Upcycled-FedProx can still converge when data is highly heterogeneous.
 Assumption 5.7.
 $||\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}||\leq h,\forall m$ and $||\nabla f(\omega)||\leq d,\forall\omega$.
-Assumption 4.7 is standard and has been used when proving the convergence of FL algorithms [34, 35]; it requires that the difference of aggregated weights between two consecutive iterations and the gradient $||\nabla f(\omega)||$ are bounded.
+Assumption 5.7 is standard and has been used when proving the convergence of FL algorithms (Li et al., 2019; Yang et al., 2022); it requires that the difference of aggregated weights between two consecutive iterations and the gradient $||\nabla f(\omega)||$ are bounded.
 Under this assumption, we have the following corollary.
 Corollary 5.8 (Convergence to the stationary point).
-Under Assumption 4.2, 4.4, and 4.7, for fixed $\mu,K$, if $\lambda_{m}$ is taken such that $\frac{\mu}{\mu+\lambda_{m}}=\mathcal{O}\Big{(}\frac{1}{\sqrt{M}}\Big{)}$, then the convergence rate of Upcycled-FL reduces to $\mathcal{O}(\frac{1}{\sqrt{M}})$.
-Corollary 4.8 provides guidance on how to choose the value of $\lambda_{m}$ to guarantee the convergence of Upcycled-FL, i.e., by taking an increasing sequence of $\{\lambda_{m}\}_{m=1}^{M}$.
+Under Assumption 5.2, 5.4, and 5.7, for fixed $\mu,K$, if $\lambda_{m}$ is taken such that $\frac{\mu}{\mu+\lambda_{m}}=\mathcal{O}\Big{(}\frac{1}{\sqrt{M}}\Big{)}$, then the convergence rate of Upcycled-FedProx reduces to $\mathcal{O}(\frac{1}{\sqrt{M}})$.
+Corollary 5.8 provides guidance on selecting the value of $\lambda_{m}$ properly to guarantee the convergence of Upcycled-FedProx, i.e., by taking an increasing sequence of $\{\lambda_{m}\}_{m=1}^{M}$.
 Intuitively, increasing $\lambda_{m}$ during the training helps stabilize the algorithm, because the deviation of local models from the previous update gets penalized more under a larger $\lambda_{m}$.
@@
 In this section, we present a privacy-preserving version of Upcycled-FL.
-Many perturbation mechanisms can be adopted to achieve differential privacy such as objective perturbation [36, 37], output perturbation [36, 38], gradient perturbation [39, 40], etc.
-Next, we will use output and objective perturbation as examples to illustrate that Upcycled-FL, by reducing the total information leakage, has a better privacy-accuracy trade-off than other methods such as FedProx, FedAvg, etc.
-Note that both output and objective perturbation methods are used for generating private updates at odd iterations, which can be used directly for even updates.
+Many perturbation mechanisms can be adopted to achieve differential privacy such as objective perturbation (Chaudhuri et al., 2011; Kifer et al., 2012), output perturbation (Chaudhuri et al., 2011; Zhang et al., 2017), gradient perturbation (Bassily et al., 2014; Wang et al., 2017), etc.
+In this section, we use output and objective perturbation as examples to illustrate that FL algorithms combined with Upcycled-FL strategy, by reducing the total information leakage, can attain a better privacy-accuracy trade-off.
+Note that both output and objective perturbation methods are used to generate private updates at odd iterations, which can be used directly for even updates.
 Output perturbation: the private odd updates $\widehat{\omega}_{i}^{2m-1}$ are generated by first clipping the local models ${\omega}_{i}^{2m-1}$ and then adding a noise random vector $n_{i}^{m}$ to the clipped model:
@@
 where parameter $\tau>0$ is the clipping threshold; the clipping ensures that if $||{\omega}_{i}^{2m-1}||_{2}\leq\tau$, then update remains the same, otherwise it is scaled to be of norm $\tau$.
-Objective perturbation: a random linear term $\langle n^{m}_{i},\omega\rangle$ is added to the objective function in odd $(2m+1)$-th iteration, and the private local model $\widehat{\omega}_{i}^{2m+1}$ is found by solving a perturbed optimization:
+Objective perturbation: a random linear term $\langle n^{m}_{i},\omega\rangle$ is added to the objective function in odd $(2m+1)$-th iteration, and the private local model $\widehat{\omega}_{i}^{2m+1}$ is found by solving a perturbed optimization.
+Taking Upcycled-FedProx as the example, we have:
 $\widehat{\omega}_{i}^{2m+1}=\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{2m}||^{2}+\langle n^{m}_{i},\omega\rangle,$
@@
 For any $m=1,2,\cdots$, if the total privacy loss up to $12m-$ 1 can be bounded by $\varepsilon_{m}$, then the total privacy loss up to the $2m$-th iteration can also be bounded by $\varepsilon_{m}$.
-Lemma 5.1 is straightforward; it can be derived directly by leveraging a property of differential privacy called immunity to post-processing [41], i.e., a differentially private output followed by any data-independent computation remains satisfying differential privacy.
+Lemma 6.1 is straightforward; it can be derived directly by leveraging a property of differential privacy called immunity to post-processing (Dwork et al., 2014), i.e., a differentially private output followed by any data-independent computation remains satisfying differential privacy.
 Based on Lemma 6.1, we can quantify the total privacy loss of private Upcycled-FL.
-We shall adopt moments accountant method [27] for output perturbation, and the analysis method in [36, 42, 43] for objective perturbation.
-These theoretical results are presented in Appendix C, Theorem C.1-C.2.
+We shall adopt moments accountant method (Abadi et al., 2016) for output perturbation, and the analysis method in (Chaudhuri et al., 2011; Zhang and Zhu, 2016; Zhang et al., 2018b) for objective perturbation.
+Here, we focus on settings where local loss function $F_{i}(\omega_{i},\mathcal{D}_{i}):=\frac{1}{|\mathcal{D}_{i}|}\sum_{d\in\mathcal{D}_{i}}\hat{F}_{i}(\omega_{i},d)$ for some $\hat{F}_{i}$, and the guarantee of privacy is presented in Theorem 6.2 (output perturbation) and 6.3 (objective perturbation) below.
+The total privacy loss in the following theorem is composed using moments accountant method (Abadi et al., 2016).
+Theorem 6.2.
+Consider the private Upcycled-FL over $2M$ iterations under output perturbation with noise $n_{i}^{m}\sim\mathcal{N}(0,\sigma^{2}\textbf{I})$, then for any $\varepsilon\geq\frac{M\tau^{2}}{2\sigma^{2}|\mathcal{D}_{i}|^{2}}$, the algorithm is $(\varepsilon,\delta)$-DP for agent $i$ for
+$\delta=\exp\left(-\frac{M\tau^{2}}{2\sigma^{2}|\mathcal{D}_{i}|^{2}}\Big{(}\frac{\varepsilon\sigma^{2}|\mathcal{D}_{i}|^{2}}{M\tau^{2}}-\frac{1}{2}\Big{)}^{2}\right).$
+Equivalently, for any $\delta\in[0,1]$, the algorithm is $(\varepsilon,\delta)$-DP for agent $i$ for
+$\varepsilon=2\sqrt{\frac{M\tau^{2}}{2\sigma^{2}|\mathcal{D}_{i}|^{2}}\log(\frac{1}{\delta})}+\frac{M\tau^{2}}{2\sigma^{2}|\mathcal{D}_{i}|^{2}}.$
+Theorem 6.3.
+Consider the private Upcycled-FL over $2M$ iterations under objective perturbation with noise $n^{m}_{i}\sim\exp{(-\alpha_{i}^{m}||n_{i}^{m}||_{2})}$.
+Suppose $\hat{F}_{i}$ is generalized linear model (Iyengar et al., 2019; Bassily et al., 2014)In supervised learning, the sample $d=(x,y)$ corresponds to the feature and label pair.
+Function $\hat{F}_{i}(\omega,d)$ is generalized linear model if it can be written as a function of $\omega^{T}x$ and $y$.
+that satisfies $||\nabla\hat{F}_{i}(\omega;d)||<u_{1},~{}|\hat{F}^{\prime\prime}_{i}|\leq u_{2}$.
+Let feature vectors be normalized such that its norm is no greater than 1, and suppose $u_{2}\leq 0.5|\mathcal{D}_{i}|\mu$ holds.
+Then the algorithm satisfies $(\varepsilon,0)$-DP for agent $i$ where $\varepsilon=\sum_{m=0}^{M}\frac{2\alpha^{m}_{i}u_{1}\mu+2.8u_{2}}{|\mathcal{D}_{i}|\mu}.$
+The assumptions on $\hat{F}_{i}$ are again fairly standard in the literature, see e.g., (Chaudhuri et al., 2011; Zhang and Zhu, 2016; Zhang et al., 2018b).
+Theorem 6.2 and 6.3 show that the total privacy loss experienced by each agent accumulates over iterations and privacy loss only comes from odd iterations.
+In contrast, if consider differentially private FedProx, accumulated privacy loss would come from all iterations.
+Therefore, to achieve the same privacy guarantee, private Upcycled-FL requires much less perturbation per iteration than private FedProx.
+As a result, accuracy can be improved significantly.
+Experiments in Section 7 show that Upcycled-FL significantly improves privacy-accuracy trade-off compared to other methods.
 7 Experiments
-In this section, we empirically evaluate Upcycled-FL and compare it with two algorithms: FedAvg [44] and FedProx [2].
+In this section, we empirically evaluate the performance of Upcycled-FL by combining it with several popular FL methods.
 We first consider non-private algorithms to examine the convergence (rate) and robustness of Upcycled-FL against statistical/device heterogeneity.
 Then, we adopt both output and objective perturbation to evaluate the private Upcycled-FL.
-Experimental setup. We run experiments on both synthetic and real data.
-  • Synthetic data: using the method in [2], we generate Synthetic(iid), Synthetic(0,0), Synthetic(0.5,0.5), Synthetic(1,1), four synthetic datasets with increasing statistical heterogeneity.
-  • Real data: we adopt two real datasets: 1) Femnist, a federated version of Emnist [45]; 2) Sentiment140(Sent140), a text sentiment analysis task on tweets [46].
-Both of the datasets can be obtained from LEAF [47].
+7.1 Datasets and Networks
+We conduct experiments on both synthetic and real data, as detailed below. More details of each dataset are given in Appendix F.1.
+Synthetic data.
+Using the method in Li et al.
+(2020), we generate Syn(iid), Syn(0,0), Syn(0.5,0.5), Syn(1,1), four synthetic datasets with increasing statistical heterogeneity.
+We use logistic regression for synthetic data.
+Real data.
+We adopt two real datasets: 1) FEMNIST, a federated version of EMNIST (Cohen et al., 2017).
+Here, A multilayer perceptron (MLP) consisting of two linear layers with a hidden dimension of 14x14, interconnected by ReLU activation functions, is used to learn from FEMNIST; 2) Sentiment140 (Sent140), a text sentiment analysis task on tweets (Go et al., 2009).
+In this context, a bidirectional LSTM with 256 hidden dimensions and 300 embedding dimensions is used to train on Sent140 dataset.
+7.2 Experimental setup
+All experiments are conducted on a server equipped with multiple NVIDIA A5000 GPUs, two AMD EPYC 7313 CPUs, and 256GB memory.
+The code is implemented with Python 3.8 and PyTorch 1.13.0 on Ubuntu 20.04.
+We employ SGD as the local optimizer, with a momentum of 0.5, and set the number of local update epochs E to 10 at each iteration $m$.
+Note that without privacy concerns, any classifier and loss function can be plugged into Upcycled-FL.
+However, if we adopt objective perturbation as privacy protection, the loss function should also satisfy assumptions in Theorem 6.3.
+We take the cross-entropy loss as our loss function throughout all experiments.
 To simulate device heterogeneity, we randomly select a fraction of devices to train at each round, and assume there are stragglers that cannot train for full rounds; both devices and stragglers are selected by random seed to ensure they are the same for all algorithms.
-We also consider several learning algorithms: we learn logistic regression on synthetic data; multilayer perceptron (MLP) and deep CNN on Femnist; Stacked LSTM on Sent140.
-For Femnist, we present the results on MLP here and other results in Appendix H.
-Because even iterations of Upcycled-FL only involve addition/subtraction operations with almost no computational cost, we train Upcycled-FL with double iterations compared to FedAvg and FedProx in approximately same training time.
-Unless explicitly stated, the results we report are averaged outcomes over all devices.
-More details of experimental setup are in Appendix H.1.
-(a) Synthetic(0.5,0.5)
-(b) Femnist
-Figure 2: Convergence of Upcycled-FL, FedProx, FedAvg under the approximate same training time (left) and impact of $\lambda_{m}$ on convergence rate of Upcycled-FL (right).
-Table 1: Training(testing) accuracy of FedAvg, FedProx, Upcycled-FL with 90% straggler: FedAvg/FedProx/Upcycled-FL are trained over synthetic data for 80/80/160 iterations, and are trained over Femnist for 150/150/300 iterations.
-For Upcycled-FL, $\lambda=0.04,0.12,0.21,0.43$ is set based on the value of $\mu=0.1,0.3,0.5,1$, respectively.
-Upcycled-FL significantly outperforms FedAvg and FedProx.
-Dataset Proximal term Accuracy(%)
-FedAvg FedProx Upcycled-FL
-Synthetic(iid) $\mu=0.1$ 97.80%(97.51%) 96.86%(96.34%) 97.88%(97.51%)
-$\mu=0.3$ 95.18%(94.88%) 96.31%(96.05%)
-$\mu=0.5$ 94.03%(93.70%) 95.40%(95.02%)
-$\mu=1$ 92.14%(91.80%) 93.76%(93.70%)
-Synthetic(0,0) $\mu=0.1$ 78.41%(76.27%) 81.66%(79.50%) 84.41%(82.44%)
-$\mu=0.3$ 81.73%(79.35%) 83.83%(81.58%)
-$\mu=0.5$ 81.55%(79.35%) 82.84%(81.29%)
-$\mu=1$ 81.22%(78.85%) 82.41%(80.57%)
-Synthetic(0.5,0.5) $\mu=0.1$ 79.53%(78.97%) 79.63%(79.58%) 81.19%(81.54%)
-$\mu=0.3$ 80.31%(80.56%) 82.14%(83.74%)
-$\mu=0.5$ 79.55%(79.46%) 82.61%(82.64%)
-$\mu=1$ 80.11%(80.20%) 82.74%(82.64%)
-Synthetic(1,1) $\mu=0.1$ 69.70%(71.12%) 79.20%(78.83%) 77.38%(78.64%)
-$\mu=0.3$ 74.77%(75.91%) 80.43%(81.47%)
-$\mu=0.5$ 75.36%(76.67%) 82.78%(83.07%)
-$\mu=1$ 75.69%(76.58%) 82.88%(82.69%)
-Femnist $\mu=0.1$ 20.43%(20.01%) 82.86%(83.73%) 86.37%(85.91%)
-$\mu=0.3$ 76.10%(76.66%) 81.94%(82.12%)
-$\mu=0.5$ 68.28%(68.87%) 77.77%(77.60%)
-$\mu=1$ 28.82%(29.05%) 54.81%(54.00%)
-Sent140 $\mu=0.1$ 73.41%(71.13%) 75.29%(72.57%) 78.12%(72.38%)
-$\mu=0.3$ 72.86%(71.38%) 74.27%(73.02%)
-$\mu=0.5$ 70.18%(69.07%) 72.92%(71.44%)
-$\mu=1$ 62.10%(62.13%) 63.91%(64.11%)
+Baselines.
+To evaluate the effectiveness of our strategy, we apply our Upcycled-FL method to seven representative methods in federated learning.
+We use the grid search to find the optimal hyperparameters for all methods to make a fair comparison.
+These methods include:
+  • FedAvg (McMahan et al., 2017): FedAvg learns the global model by averaging the client’s local model.
+  • FedAvgM (Hsu et al., 2019): FedAvgM introduces momentum while averaging local models to improve convergence rates and model performance, especially in non-i.i.d. settings.
+  • FedProx (Li et al., 2020): FedProx adds a proximal term to the local objective function, which enables more robust convergence when data is non-i.i.d. across different clients.
+  • Scaffold (Karimireddy et al., 2020): Scaffold uses control variates to correct the local updates, which helps in dealing with non-i.i.d. data and accelerates convergence.
+  • FedDyn (Acar et al., 2021): FedDyn considers a dynamic regularization term to align the local model updates more closely with the global model.
+  • pFedMe (T Dinh et al., 2020): pFedMe is a personalization method to handle client heterogeneity. We set the hyperparameter $k$ in pFedMe as 5 to accelerate the training.
+  • FedYogi (Reddi et al., 2021): FedYogi considers the adaptive optimization for the global model aggregation.
+Unless explicitly stated, the results we report are averaged outcomes over all devices. More details of experimental setup are in Appendix F.1.
+Table 1: Average accuracy and standard deviation with 90% straggler on the testing dataset in the non-private setting over four runs: models are trained over synthetic data (Syn) for 80 iterations (160 for upcycled version), trained over Femnist for 150 iterations (300 for upcycled version), and trained over Sent140 for 80 iterations (160 for upcycled version).
+We use the grid search to find the optimal results for all methods.
+Method Dataset
+Syn(iid) Syn(0,0) Syn(0.5,0.5) Syn(1,1) FEMNIST Sent140
+FedAvg 98.06±0.07 79.28±0.61 81.58±0.43 80.40±1.28 81.38±3.54 76.11±0.11
+Upcycled-FedAvg 98.83±0.29 81.46±0.48 82.89±0.17 81.49±0.53 82.10±1.11 76.32±0.45
+FedAvgM 98.43±0.07 80.29±0.83 82.60±0.37 80.59±1.28 80.15±3.72 75.7±0.85
+Upcycled-FedAvgM 98.72±0.47 81.74±0.42 83.13±0.12 81.37±0.82 81.30±5.23 74.88±2.29
+FedProx 96.52±0.07 80.72±0.77 81.99±0.55 81.19±0.19 79.35±0.65 73.94±0.13
+Upcycled-FedProx 97.62±0.32 80.88±0.97 83.10±0.83 81.94±0.57 80.33±3.43 74.25±0.34
+Scaffold 97.51±0.24 80.26±1.54 82.44±1.66 74.91±2.67 76.83±2.97 76.34±0.56
+Upcycled-Scaffold 98.68±0.12 81.10±0.57 82.64±1.39 76.14±1.28 77.88±5.36 77.34±0.22
+FedDyn 97.00±0.19 81.62±0.97 80.64±0.81 77.27±2.95 81.76±0.98 75.97±0.35
+Upcycled-FedDyn 98.32±0.08 82.41±1.04 82.89±0.80 80.03±3.02 83.33±0.71 76.03±0.58
+pFedme 96.30±0.14 89.15±0.24 89.43±0.67 93.06±0.27 71.73±4.30 72.81±0.85
+Upcycled-pFedme 96.77±0.12 89.08±0.22 89.55±0.62 93.12±0.23 76.66±3.37 74.07±0.74
+FedYogi $99.30_{\pm 0.35}$ $81.20_{\pm 2.64}$ $79.49_{\pm 1.30}$ $78.95_{\pm 1.98}$ $73.53_{\pm 7.73}$ $77.02_{\pm 0.07}$
+Upcycled-FedYogi $\textbf{99.41}_{\pm 0.32}$ $\textbf{81.65}_{\pm 2.44}$ $\textbf{80.84}_{\pm 1.30}$ $\textbf{80.17}_{\pm 0.84}$ $\textbf{75.64}_{\pm 3.17}$ $\textbf{77.59}_{\pm 0.26}$
+7.3 Results
 Convergence and Heterogeneity.
-We evaluate the convergence rate, training time, and accuracy of Upcycled-FL under different parameter settings.
-We take $\lambda_{m}=\lambda,\forall m$ in experiments, while we observe similar results for time-varying $\lambda_{m}$.
-Table 1 compares the training/testing accuracy of different algorithms when the device heterogeneity is high (90% stragglers).
-The results show that Upcycled-FL outperforms both FedProx and FedAvg on all datasets (with appropriate $\mu$).
-While FedAvg achieves good performance on Synthetic(iid), it is not robust to heterogeneous data.
-When data is i.i.d., adding the proximal term $\frac{\mu}{2}||\omega-\overline{\omega}^{t}||^{2}$ may hurt the performance.
-However, the proximal term can help stabilize the algorithm and significantly improve the performance in practical settings when data is heterogeneous; these observations are consistent with [2].
-Importantly, Upcycled-FL is more robust to the level of statistical heterogeneity than FedProx as it could attain consistent improvements for all settings.
-In practice, $\mu$ needs to be carefully tuned; with a carefully selected $\mu$, Upcycled-FL could achieve $2\%\sim 7\%$ improvement on accuracy compared to FedProx.
-Figure 2(a) and 2(b) illustrate the convergence property of Upcycled-FL on Synthetic(0.5,0.5) and Femnist (in comparison with FedProx and FedAvg), and examine the impact of hyper-parameter $\lambda_{m}$.
-Loss in $y$-axes indicates the averaged loss over all devices.
+Because even iterations of Upcycled-FL only involve addition/subtraction operations with no transmission overhead and almost no computational cost, we train the Upcycled version of FL algorithms with double iterations compared to baselines in approximately same training time in this experiment.
+We evaluate the convergence rate and accuracy of Upcycled-FL under different dataset and heterogeneity settings.
 In each iteration, 30% of devices are selected with 90% stragglers.
-We observe that Upcycled-FL could attain a comparable performance as FedProx in the first 80 (resp.
-240) iterations on Synthetic(0.5,0.5) (resp.
-Femnist), with 50% (resp.
-20%) of SGD computation reduced.
-Specifically, the final training (testing) accuracy attained on Synthetic(0.5,0.5) is 82.61% (82.64%) for Upcycled-FL, 79.55% (79.46%) for FedProx, and 79.53% (78.97%) for FedAvg; while the accuracy attained on Femnist is 81.94% (82.12%) for Upcycled-FL, 76.10% (76.66%) for FedProx, and 20.43% (20.01%) for FedAvg.
-More results on other datasets are presented in Appendix H.4.
-(a) Synthetic(iid)
-(b) Synthetic(0,0)
-(c) Synthetic(iid)
-(d) Synthetic(0,0)
-Figure 3: Comparison of private Upcycled-FL and private FedProx using objective perturbation ((a)(b)) and output perturbation ((c)(d)) with $\delta=0.001$.
-We take $\mu=0.1,\lambda=0.04$ for all datasets.
-$\bar{\epsilon}$ denotes the average $\epsilon$ among all devices.
-In both experiments, the privacy of private Upcycled-FL is strictly stronger than private FedProx.
-The effect of hyper-parameter $\lambda_{m}=\lambda$ on the convergence rate of Upcycled-FL is also evaluated in Figure 2(a)2(b) (right).
-Although the sufficient conditions in Theorem 4.6 and Corollary 4.8 suggest that Upcycled-FL converges when $\lambda$ is sufficiently large, the experiments indicate that Upcycled-FL can easily converge in practice even with small $\lambda$.
-Furthermore, we observe that models with smaller $\lambda$ tend to converge faster.
-This is because that $\lambda$ controls step size $\frac{\mu}{\mu+\lambda}$ (by Eqn.
-(5)): a larger $\lambda$ results in a smaller step size and thus slower convergence.
-Privacy-Accuracy Tradeoff.
-We next inspect the accuracy-privacy trade-off of private Upcycled-FL and compare it with private FedProx.
-We adopt both objective and output perturbation to preserve privacy in the experiments, while other privacy-preserving techniques can also be used.
-For each parameter setting, we perform 10 independent runs of the experiment and record both the mean and range of their losses.
-To precisely quantify privacy, we focus on settings without device heterogeneity (no straggler) and low statistical heterogeneity (Synthetic(iid), Synthetic(0,0)).
-Figure 3(a) and 3(b) demonstrate the performance (averaged loss and the respective privacy loss $\epsilon$) of (non-)private Upcylced-FL and private FedProx on synthetic data using objective perturbation.
-As expected, private Upcycled-FL is much more stable and significantly improves both privacy and accuracy over private FedProx.
-This is because the private Upcycled-FL with less information leakage requires much less perturbation to attain the same privacy guarantee as private FedProx.
-In general, we observe that private Upcycled-FL can experience 9.5% less privacy loss than private FedProx, and meanwhile having 11.0% (or 11.5%) higher training (or testing) accuracy on average and reducing 48% of training time.
-We also conduct similar experiments using output perturbation (Figure 3(c) and 3(d)).
-Results show that on average, private Upcycled-FL can experience 12.0% less privacy loss than private FedProx while having 6.1% (or 6.2%) higher training (or testing) accuracy; the training time is still reduced by 48% under Upcycled-FL.
-We also conduct experiments to examine the accuracy-privacy trade-off under different magnitudes of perturbation, the results are presented in Appendix H.5.1.
-It shows that private Upcycled-FL can consistently outperform private FedProx under different levels of privacy.
+Table 1 compares the average accuracy of different algorithms when the device heterogeneity is high (90% stragglers).
+The results show that all baselines can be enhanced by our methods.
+Notably, while FedAvg achieves good performance on Syn(iid), it is not robust on heterogeneous data, e.g.
+Syn(0,0), Syn(0.5,0.5), and Syn(1,1).
+Nonetheless, Upcycled-FedAvg makes it comparable with the regular FedProx algorithm, which shows that our strategy can also mitigate performance deterioration induced by data heterogeneity.
+When data is i.i.d., FedProx with the proximal term $\frac{\mu}{2}||\omega-\overline{\omega}^{t}||^{2}$ may hurt the performance compared with FedAvg.
+However, the proximal term can help stabilize the algorithm and significantly improve the performance in practical settings when data is heterogeneous; these observations are consistent with (Li et al., 2020).
+Importantly, Upcycled-FL strategy further makes it more robust to statistical heterogeneity than regular FedProx as it could attain consistent improvements for all settings.
+(a) Syn(iid)
+(b) Syn(0.5,0.5)
+(c) FEMNIST
+Figure 2: Comparison on average loss and standard deviation between Upcycled-FL methods and original FL algorithms in the non-private setting under the approximate same training time.
+The training time refers to the time needed for a given number of iterations.
+Upcycled-FL does not require an update in the even iterations, allowing Upcycled-FL to train with doubled iterations.
+(a) Syn(iid)
+(b) Syn(0,0)
+(c) Syn(0.5,0.5)
+(d) Syn(1,1)
+Figure 3: Comparison on average loss and standard deviation of private Upcycled-FL and private FL methods using output perturbation.
+The noise parameter $\sigma$ is 1.0 for all baselines, while $\sigma$ of the Upcycled version is set to 0.8.
+Taking the iid dataset as an example, $\bar{\epsilon}=1.40$ for the Upcycled version, which ensures stronger privacy than the original method with $\bar{\epsilon}=1.59$.
+Figure 2(a), 2(b) and 2(c) compare the convergence property on Syn(iid), Syn(0.5,0.5) and FEMNIST.
+Note that, by using the aggregation rule in Figure 1(b), the training time of Upcycled-FL is almost the same as the baselines without introducing extra cost.
+We observe that under the same training time (the number of iterations for baselines), Upcycled-FL strategy brings benefits (achieving lower loss along training) for baselines in most cases.
+And this improvement is significant when the dataset is i.i.d.
+The loss trends are consistent with results in Table 1.
+We provide more results on the other three datasets in Appendix F.3.
+(a) Syn(iid)
+(b) Syn(0,0)
+(c) Syn(0.5,0.5)
+(d) Syn(1,1)
+Figure 4: Comparison on average loss and standard deviation of private Upcycled-FL and private FL methods using objective perturbation.
+Under objective perturbation, noise parameter $\alpha$ is 10 for all baselines, while $\alpha$ of the Upcycled version is set to 20 to ensure stronger privacy than the original versions.
+Taking the iid dataset as an example, $\bar{\epsilon}$ associated with these noise parameters is 7.36 for FedProx and 7.25 for Upcycled-FedProx (when $\mu$ = 0.5).
+Privacy-Accuracy Trade-off.
+We next inspect the accuracy-privacy trade-off of private Upcycled-FL and compare it with private baselines.
+Although we adopt both objective and output perturbation to achieve differential privacy, other techniques can also be used.
+For each parameter setting, we still conduct a grid search and perform 4 independent runs of experiments.
+To precisely quantify privacy, no straggler is considered in this experiment.
+We reported the results using output perturbation and objective perturbation.
+Figure 3 and 4 demonstrate the performance of private Upcylced-FL and private baselines on synthetic data using two types of perturbation respectively.
+We also report the results on real data in Appendix F.4.
+Here we carefully set the perturbation strength of each algorithm such that the privacy loss $\epsilon$ for private Upcycled-FL is strictly less than the original methods.
+As expected, private Upcycled-FL is more stable than baselines and attains a lower loss value under smaller $\epsilon$.
+This is because the private Upcycled-FL with less information leakage requires much less perturbation to attain the same privacy guarantee as FL methods under privacy constraints.
+We also observe that in general Upcycled-FL can be used to augment all baseline methods with or without client local heterogeneity.
 8 Conclusion
-This paper proposes Upcycled-FL, a novel federated learning framework under which information leakage and computation cost can be reduced significantly.
-We theoretically examined the convergence (rate) of Upcycled-FL.
-The experiments show that compared to existing algorithms, Upcycled-FL can be more robust on heterogeneous data and attain much better privacy-accuracy trade-off under common differential privacy mechanisms.
+This paper proposes Upcycled-FL, a novel plug-in federated learning strategy under which information leakage and computation costs can be reduced significantly.
+We theoretically examined the convergence (rate) of Upcycled-FedProx, a special case when Upcycled-FL strategy is applied to a well-known FL algorithm named FedProx.
+Extensive experiments on synthetic and read data further show that Upcycled-FL can be combined with common FL algorithms and enhance their robustness on heterogeneous data while attaining much better privacy-accuracy trade-off under common differential privacy mechanisms.
+Acknowledgments
+This material is based upon work supported by the U.S.
+National Science Foundation under awards IIS-2040800, IIS-2112471, IIS-2202699, IIS-2301599, and CMMI-2301601, by grants from the Ohio State University’s Translational Data Analytics Institute and College of Engineering Strategic Research Initiative.
 Appendix A Notation Table
@@
 $n_{i}^{t}$ random noise added to agent $i$ at time $t$
-$\mu$ hyper-parameter for proximal term in FedProx and Upcycled-FL
+$\mu$ hyper-parameter for proximal term in FedProx and Upcycled FedProx
 $\lambda_{m}$ hyper-parameter for first-order approximation at even iteration $2m$ in Upcycled-FL
-Appendix B Related Work
-Differentially private federated learning.
-Differential privacy has been widely used in federated learning to provide privacy guarantees [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20].
-For example, [7] uses the Gaussian mechanism for a federated learning problem and propose an incentive mechanism to encourage users to share their data and participate in the training process.
-[8] introduces $f$-differential privacy, a generalized version of Gaussian differential privacy, and propose a federated learning algorithm satisfying this new notion.
-[9] proposes a new mechanism called Random Response with Priori (RRP) to achieve local differential privacy and apply this mechanism to the text data by training a Latent Dirichlet Allocation (LDA) model using a federated learning algorithm.
-[20] adapts the Bayesian privacy accounting method to the federated setting and propose joint accounting method for estimating client-level and instance-level privacy simultaneously and securely.
-[19] presents a private scheme that adds noise to parameters at the random selected devices before aggregating and provides a convergence bound.
-[10] combines the Gaussian mechanism with gradient clipping in federated learning to improve the privacy-accuracy tradeoff.
-[5] considers a different setting where only the last update is publicly released and the central server and other devices are assumed to be trustworthy.
-However, all these algorithms need client’s local data to update the model and the information leakage inevitably occurs at every iteration.
-This is fundamentally different from Upcycled-FL, which reuses existing results to update half of iterations and significantly reduces information leakage and computation.
-Tackling heterogeneity in federated learning.
-It’s worth mentioning that, empirically Upcycled-FL also outperforms existing baseline algorithm under device and statistical heterogeneity.
-In real-world scenarios, local data are often non-identically distributed across different devices; different devices are also often equipped with different specifications and computation capabilities.
-Such heterogeneity often causes instability in the model performance and leads to divergence.
-Many approaches have been proposed to tackle this issue.
-For example, FedAvg [44] uses a random selection of devices at each iteration to reduce the negative impact of statistical heterogeneity; however, it may fail to converge when heterogeneity increases.
-Other methods include FedProx [2], a generalization and re-parameterization of FedAvg that adds a proximal term to the objective function to penalize deviations in the local model from the previous aggregation, and FedNova [3] that re-normalizes local updates before updating to eliminate objective inconsistency.
-It turns out that Upcycled-FL exhibits superior performance in the presence of heterogeneity because gradients encapsulate information on data heterogeneity, the reusing of which leads to a boost in performance.
-Appendix C Privacy analysis and results
-We provide the privacy analysis of private Upcycled-FL under output and objective perturbation.
-Here, we focus on settings where local loss function $F_{i}(\omega_{i},\mathcal{D}_{i}):=\frac{1}{|\mathcal{D}_{i}|}\sum_{d\in\mathcal{D}_{i}}\hat{F}_{i}(\omega_{i},d)$ for some $\hat{F}_{i}$, and the guarantee of privacy is presented in Theorem C.1 (output perturbation) and C.2 (objective perturbation) below.
-The total privacy loss in the following theorem is composed using moments accountant method [27].
-Theorem C.1.
-Consider the private Upcycled-FL over $2M$ iterations under output perturbation with noise $n_{i}^{m}\sim\mathcal{N}(0,\sigma^{2}\textbf{I})$, then for any $\varepsilon\geq\frac{M\tau^{2}}{2\sigma^{2}|\mathcal{D}_{i}|^{2}}$, the algorithm is $(\varepsilon,\delta)$-DP for agent $i$ for
-$\delta=\exp\left(-\frac{M\tau^{2}}{2\sigma^{2}|\mathcal{D}_{i}|^{2}}\Big{(}\frac{\varepsilon\sigma^{2}|\mathcal{D}_{i}|^{2}}{M\tau^{2}}-\frac{1}{2}\Big{)}^{2}\right).$
-Equivalently, for any $\delta\in[0,1]$, the algorithm is $(\varepsilon,\delta)$-DP for agent $i$ for
-$\varepsilon=2\sqrt{\frac{M\tau^{2}}{2\sigma^{2}|\mathcal{D}_{i}|^{2}}\log(\frac{1}{\delta})}+\frac{M\tau^{2}}{2\sigma^{2}|\mathcal{D}_{i}|^{2}}.$
-Theorem C.2.
-Consider the private Upcycled-FL over $2M$ iterations under objective perturbation with noise $n^{m}_{i}\sim\exp{(-\alpha_{i}^{m}||n_{i}^{m}||_{2})}$.
-Suppose $\hat{F}_{i}$ is generalized linear model [48, 39]In supervised learning, the sample $d=(x,y)$ corresponds to the feature and label pair.
-Function $\hat{F}_{i}(\omega,d)$ is generalized linear model if it can be written as a function of $\omega^{T}x$ and $y$.
-that satisfies $||\nabla\hat{F}_{i}(\omega;d)||<u_{1},~{}|\hat{F}^{\prime\prime}_{i}|\leq u_{2}$.
-Let feature vectors be normalized such that its norm is no greater than 1, and suppose $u_{2}\leq 0.5|\mathcal{D}_{i}|\mu$ holds.
-Then the algorithm satisfies $(\varepsilon,0)$-DP for agent $i$ where $\varepsilon=\sum_{m=0}^{M}\frac{2\alpha^{m}_{i}u_{1}\mu+2.8u_{2}}{|\mathcal{D}_{i}|\mu}.$
-The assumptions on $\hat{F}_{i}$ are again fairly standard in the literature, see e.g., [36, 42, 43].
-Theorem C.1 and C.2 show that the total privacy loss experienced by each agent accumulates over iterations and privacy loss only comes from odd iterations.
-In contrast, if consider differentially private FedProx, accumulated privacy loss would come from all iterations.
-Therefore, to achieve the same privacy guarantee, private Upcycled-FL requires much less perturbation per iteration than private FedProx.
-As a result, accuracy can be improved significantly.
-Experiments in Section 6 show that Upcycled-FL significantly improves privacy-accuracy trade-off compared to other methods.
-Appendix D Lemmas
-Lemma D.1.
-Define $\widetilde{\omega}^{t}=\mathbb{E}_{i}(\omega_{i}^{t})$. Suppose conditions in Theorem 4.6 hold. Then the following holds for all $m$:
+$\tau$ the clipping threshold for output perturbation
+Appendix B Lemmas
+Lemma B.1.
+Define $\widetilde{\omega}^{t}=\mathbb{E}_{i}(\omega_{i}^{t})$. Suppose conditions in Theorem 5.6 hold, then the following holds $\forall m$:
 $f(\widetilde{\omega}^{2m+1})$$\leq$ $f(\overline{\omega}^{2m-1})-\widehat{C}_{1}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho}\Big{)}||\nabla f(\overline{\omega}^{2m-1})||^{2}$
@@
 $\widehat{C}_{3}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{\mu}{\mu+\lambda_{m}}\Big{)}$ $=$ $\frac{L(L+\rho)^{2}}{2\rho^{2}}\frac{\mu^{2}}{(\mu+\lambda_{m})^{2}}$
-Lemma D.2.
-Let $\mathcal{S}_{m}$ be the set of $K$ randomly selected local devices got updated at iterations $12m-$ 1 and $2m$, and $\mathbb{E}_{\mathcal{S}_{m}}[\cdot]$ be expectation with respect to the choice of devices.
+Lemma B.2.
+Let $\mathcal{S}_{m}$ be the set of $K$ randomly selected local devices got updated at iterations $12m-$ 1 and $2m$, and $\mathbb{E}_{\mathcal{S}_{m}}[\cdot]$ be the expectation with respect to the choice of devices.
 Then we have
@@
 $\widetilde{C}_{3}\Big{(}B,L,\frac{1}{K},\frac{1}{\rho},\frac{\mu}{\mu+\lambda_{m}}\Big{)}$ $=$ $\Big{(}\frac{2}{K}\frac{L^{2}}{\rho^{2}}+2L\frac{L+\rho}{\rho}\sqrt{\frac{2}{K}}\frac{L}{\rho}\Big{)}\cdot\Big{(}\frac{\mu}{\mu+\lambda_{m}}\Big{)}^{2}$
-Appendix E Proofs
-Proof of Lemma D.1
+Appendix C Proofs
+Proof of Lemma B.1
 Proof.
@@
 $\leq$ $\frac{L}{\rho}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||+\frac{1}{\rho}||\nabla F_{i}(\overline{\omega}^{2m-1})||$
-Plug in Eqn. (E),
+Plug in Eqn. (C),
 $||\Phi^{2m+1}||$$\leq$ $\frac{L}{\mu\rho}\mathbb{E}_{i}[||\nabla F_{i}(\overline{\omega}^{2m-1})||]+\Big{(}\frac{L^{2}}{\mu\rho}+\frac{L+\mu}{\mu}\Big{)}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||$
@@
 $+\frac{L(L+\rho)^{2}}{2\rho^{2}}\frac{\mu^{2}}{(\mu+\lambda_{m})^{2}}||\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}||^{2}$
-The Lemma D.1 is proved. ∎
-Proof of Lemma D.2
+The Lemma B.1 is proved. ∎
+Proof of Lemma B.2
 Proof.
@@
 Therefore,
-$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]$
-$\leq f(\widetilde{\omega}^{2m+1})$
+$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]\leq$ $f(\widetilde{\omega}^{2m+1})$
 $+\mathbb{E}_{\mathcal{S}_{m}}\Big{[}\Big{(}||\nabla f(\overline{\omega}^{2m-1})||+L(||\widetilde{\omega}^{2m+1}-\overline{\omega}^{2m-1}||+||\overline{\omega}^{2m+1}-\overline{\omega}^{2m-1}||)\Big{)}||\overline{\omega}^{2m+1}-\widetilde{\omega}^{2m+1}||\Big{]}$
@@
 $=$ $\sqrt{\frac{2}{K}}\Big{(}\frac{L}{\rho}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||+\frac{B}{\rho}||\nabla f(\overline{\omega}^{2m-1})||\Big{)}$
-By Eqn. (E),
+By Eqn. (C),
 $||\nabla f(\overline{\omega}^{2m-1})||+2L||\widetilde{\omega}^{2m+1}-\overline{\omega}^{2m-1}||\leq 2L\frac{L+\rho}{\rho}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||+\frac{2LB+\rho}{\rho}||\nabla f(\overline{\omega}^{2m-1})||$
 Re-organize, we have
-$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]$
-$\leq f(\widetilde{\omega}^{2m+1})+\frac{2}{K}\frac{L^{2}}{\rho^{2}}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||^{2}+\frac{2B^{2}}{K\rho^{2}}||\nabla f(\overline{\omega}^{2m-1})||^{2}$
+$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]\leq$ $f(\widetilde{\omega}^{2m+1})+\frac{2}{K}\frac{L^{2}}{\rho^{2}}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||^{2}+\frac{2B^{2}}{K\rho^{2}}||\nabla f(\overline{\omega}^{2m-1})||^{2}$
 $+\frac{4LB}{K\rho^{2}}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||\cdot||\nabla f(\overline{\omega}^{2m-1})||$
@@
 $+\Big{(}\frac{2B^{2}}{K\rho^{2}}+\frac{2LB+\rho}{\rho}\sqrt{\frac{2}{K}}\frac{B}{\rho}\Big{)}||\nabla f(\overline{\omega}^{2m-1})||^{2}$
-Lemma D.2 is proved. ∎
+Lemma B.2 is proved. ∎
 Proof of Lemma 5.5
 Proof.
-Lemma 4.5 can be proved by combing Lemmas D.1 and D.2, where
+Lemma 5.5 can be proved by combing Lemmas B.1 and B.2, where
 $\boldsymbol{C_{1}}$ ::=:= $C_{1}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{1}{K}\Big{)}=\widehat{C}_{1}\Big{(}L,B,\frac{1}{\mu},\frac{1}{\rho}\Big{)}-\widetilde{C}_{1}\Big{(}L,B,\frac{1}{K},\frac{1}{\rho}\Big{)}$ (11)
@@
 $=$ $\Big{(}\frac{L(L+\rho)^{2}}{2\rho^{2}}+\frac{2}{K}\frac{L^{2}}{\rho^{2}}+2L\frac{L+\rho}{\rho}\sqrt{\frac{2}{K}}\frac{L}{\rho}\Big{)}\cdot\Big{(}\frac{\mu}{\mu+\lambda_{m}}\Big{)}^{2}$
-Appendix F Proof of Theorem 4.6
+Appendix D Proof of Theorem 5.6
 This can be proved directed using Lemma 5.5 by averaging over all $M$ odd iterations.
-Appendix G Proof of Theorem C.1
+Appendix E Proof of Theorem 6.2
 WLOG, consider the case when local device got updated in every iteration and the algorithm runs over $2M$ iterations in total.
@@
 To simplify the notations, we will drop the index $i$ as we are only concerned with one agent.
-According to [27], for a mechanism $\mathscr{M}$ outputs $o$, with inputs $d$ and $\hat{d}$, let a random variable $c(o;\mathscr{M},d,\hat{d})=\log\frac{\text{Pr}(\mathscr{M}(d)=o)}{\text{Pr}(\mathscr{M}(\hat{d})=o)}$ denote the privacy loss at $o$, and
+According to Abadi et al.
+[2016], for a mechanism $\mathscr{M}$ outputs $o$, with inputs $d$ and $\hat{d}$, let a random variable $c(o;\mathscr{M},d,\hat{d})=\log\frac{\text{Pr}(\mathscr{M}(d)=o)}{\text{Pr}(\mathscr{M}(\hat{d})=o)}$ denote the privacy loss at $o$, and
 $\alpha_{\mathscr{M}}(\lambda)=\underset{d,\hat{d}}{\max}\log\mathbb{E}_{o\sim\mathscr{M}(d)}\{\exp(\lambda c(o;\mathscr{M},d,\hat{d}))\}$
@@
 $\alpha_{\{\mathscr{M}^{m}\}_{m=1}^{M}}(\lambda)$$\leq$ $\sum_{m=1}^{M}\alpha_{\mathscr{M}^{m}}(\lambda)=\frac{M\tau^{2}\lambda(\lambda+1)}{2|D|^{2}\sigma^{2}}$
-Use the tail bound [Theorem 2, [27]], for any $\varepsilon\geq\frac{M\tau^{2}}{2|D|^{2}\sigma^{2}}$, the algorithm is $(\varepsilon,\delta)$-differentially private for
+Use the tail bound [Theorem 2, Abadi et al. [2016]] , for any $\varepsilon\geq\frac{M\tau^{2}}{2|D|^{2}\sigma^{2}}$, the algorithm is $(\varepsilon,\delta)$-differentially private for
 $\delta$ $=$ $\underset{\lambda:\lambda\geq 0}{\min}~{}h(\lambda)=\underset{\lambda:\lambda\geq 0}{\min}\exp\Big{(}\frac{M\tau^{2}\lambda(\lambda+1)}{2|D|^{2}\sigma^{2}}-\lambda\varepsilon\Big{)}$
 To find $\lambda^{*}=\underset{\lambda:\lambda\geq 0}{\text{argmin}}~{}h(\lambda)$, take derivative of $h(\lambda)$ and assign 0 gives the solution $\bar{\lambda}=\frac{\varepsilon|D|^{2}\sigma^{2}}{M\tau^{2}}-\frac{1}{2}\geq 0$, and $h^{\prime\prime}(\bar{\lambda})>0$, implies $\lambda^{*}=\bar{\lambda}$.
-Plug into (G) gives:
+Plug into (E) gives:
 $\delta=\exp\left(\Big{(}\frac{M\tau^{2}}{4|D|^{2}\sigma^{2}}-\frac{\varepsilon}{2}\Big{)}\Big{(}\frac{\varepsilon|D|^{2}\sigma^{2}}{M\tau^{2}}-\frac{1}{2}\Big{)}\right)$ (15)
@@
 $\varepsilon$ $=$ $\underset{\lambda:\lambda\geq 0}{\min}~{}h_{1}(\lambda)=\underset{\lambda:\lambda\geq 0}{\min}~{}\frac{M\tau^{2}(\lambda+1)}{2|D|^{2}\sigma^{2}}+\frac{1}{\lambda}\log\left(\frac{1}{\delta}\right)=2\sqrt{\frac{M\tau^{2}}{2|D|^{2}\sigma^{2}}\log(\frac{1}{\delta})}+\frac{M\tau^{2}}{2|D|^{2}\sigma^{2}}$
-Proof of Theorem C.2
+Proof of Theorem 6.3
 Proof.
@@
 $=$ $\exp\Big{(}\sum_{m=0}^{M}\frac{2\alpha^{m}u_{1}\mu+2.8u_{2}}{|\mathcal{D}|\mu}\Big{)}$
-Theorem C.1 is proved. ∎
-Appendix H Experiments
-H.1 Details of Datasets
+Theorem 6.2 is proved. ∎
+Appendix F Experiments
+F.1 Details of Datasets
 Table 2: Details of datasets. Numbers in parentheses represent the amount of test data. All of the numbers round to integer.
@@
 mean stdev
-Synthetic iid 6726(683) 30 224 166
+Syn iid 6726(683) 30 224 166
 0,0 13791(1395) 30 460 841
@@
 Synthetic.
-The synthetic data is generated using the same method in [2].
+The synthetic data is generated using the same method in Li et al.
+[2020].
 We briefly describe the generating steps here.
@@
 We generate 4 datasets in total.
-They’re Synthetic(iid) Synthetic(0,0) with $\beta=0$ and $\gamma=0$, Synthetic(0.5,0.5) with $\beta=0.5$ and $\gamma=0.5$ and Synthetic(1,1) with $\beta=1$ and $\gamma=1$.
-Femnist: Similar with [2], we subsample 10 lower case characters (‘a’-‘j’) from EMNIST [45] and distribute 5 classes to each device. There are 50 devices in total. The input is 28*28 image.
-Sent140: A text sentiment analysis task on tweets [46]. The input is sequence of length 25 and output is the probabilities of 2 classes.
+They’re Syn(iid) Syn(0,0) with $\beta=0$ and $\gamma=0$, Syn(0.5,0.5) with $\beta=0.5$ and $\gamma=0.5$ and Syn(1,1) with $\beta=1$ and $\gamma=1$.
+In the output perturbation experiments, we set $\sigma$ to $1.0$ for the baseline methods and to $0.8$ for the Upcycled baselines, ensuring that the privacy budget $\epsilon$ for the baselines is always greater than that for the Upcycled baselines (e.g., the baseline $\bar{\epsilon}=0.773$ and the Upcycled baseline $\bar{\epsilon}=0.683$ for Syn(iid)).
+In the objective perturbation experiments, we set $\alpha$ to $10$ for the baselines and $20$ for the Upcycled baselines to achieve similar levels of information leakage, while still maintaining that $\epsilon$ for the Upcycled baselines is less than for the others.
+This constraint is maintained across all experiments.
+FEMNIST: Similar with Li et al.
+[2020], we subsample 10 lower case characters (‘a’-‘j’) from EMNIST Cohen et al.
+[2017] and distribute 5 classes to each device.
+There are 50 devices in total.
+The input is 28x28 image.
+In the privacy experiments, $\sigma$ is set to $0.27$ for the baselines and $0.2$ for the Upcycled baselines, and $\alpha$ is set to $100$ for the baselines and $200$ for the Upcycled baselines, respectively.
+Sent140: A text sentiment analysis task on tweets Go et al.
+[2009].
+The input is a sequence of length 25 and the output is the probabilities of 2 classes.
+Here, $\sigma$ is set to $0.27$ for the baselines and $0.2$ for the Upcycled baselines, and $\alpha$ is set to $15$ and $30$ respectively.
 A brief summary of dataset can be found in Table 2.
-H.2 Details of the Model
-Classifier and Loss function.
-In our experiment, we use logistic regression for Synthetic, Multilayer perceptron and CNN for Femnist, Stacked LSTM for Sent140.
-Note that without privacy concern, any classifier and loss function can be plugged in to Upcycled-FL.
-However, if we adopt objective perturbation as privacy protection, the loss function should also satisfy assumptions in Theorem C.2.
-Thus log loss $\mathscr{L}(z)=log(1+\exp(-z))$ is used.
-It’s not hard to verify that this loss function satisfies assumptions in Theorem C.2 by taking first and second order derivatives($\left|\mathscr{L}^{\prime}\right|\leq u_{1}=1\text{ and }\mathscr{L}^{\prime\prime}\leq u_{2}=\frac{1}{4}$).
-For Multilayer perceptron, we use a two layer network with hidden dimension 14*14.
-For CNN, a 3*3 convolutional layer with 3 channels is adopted, following by two linear layers(followed by activation and dropout) with dimension 1000 and 100.
-For Stacked LTSM, we use 2 layers with 128 dimension Sent140.
-The word embedding dimension of Sent140 is 300.
-Implementation and Environment.
-Our code is implemented in Pytorch 1.11 [49].
-We employ SGD as local optimizer, with momentum 0.5, and set number of epochs E to 20 at each iteration $m$.
-Learning rate is tuned to 0.05.
-We run the model on a single GTX 1080 Ti.
-H.3 CNN on Femnist
-Figure 4: Loss and accuracy on Femnist using CNN. Trained with fraction=0.3, straggler=90%. $\mu=0.1,\lambda=0.04$. Training(testing) accuracy is Upcycled-FL 96.36%(96.21%), FedProx 95.63%(95.74%).
-H.4 Convergence on all datasets
-H.4.1 90% Straggler
-Figure 5: Loss and accuracy on synthetic datasets.
-Trained with fraction=0.3, straggler=90%.
-(a) Synthetic(iid); (b) Synthetic(0,0); (c) Synthetic(0.5,0.5); (d) Synthetic(1,1).
-$\mu=0.1,\lambda=0.04$ for Synthetic(iid), Synthetic(0,0), $\mu=1,\lambda=0.42$ for Synthetic(0.5,0.5), Synthetic(1,1).
-(a) $\mu=0.1$
-(b) $\mu=0.3$
-(c) $\mu=0.5$
-Figure 6: Loss and accuracy on Femnist. Trained with fraction=0.3, straggler=90%. $\lambda=0.04,0.12,0.21$ is set based on the value of $\mu=0.1,0.3,0.5$, respectively.
-(a) $\mu=0.1$
-(b) $\mu=0.3$
-Figure 7: Loss and accuracy on Sent140. Trained with fraction=0.3, straggler=90%. $\mu=0.1,\lambda=0.04$ for (a), $\mu=0.3,\lambda=0.12$ for (b).
-H.4.2 30% Straggler
-Figure 8: Loss and accuracy on synthetic datasets.
-Trained with fraction=0.3, straggler=30%.
-(a) Synthetic(iid); (b) Synthetic(0,0); (c) Synthetic(0.5,0.5); (d) Synthetic(1,1).
-$\mu=0.1,\lambda=0.04$ for Synthetic(iid), Synthetic(0,0), $\mu=1,\lambda=0.42$ for Synthetic(0.5,0.5), Synthetic(1,1).
-H.5 Privacy on all datasets
-H.5.1 Objective perturbation
-  • Convergence
-    Figure 9: Objective perturbation on synthetic datasets.
-Trained with fraction=1, straggler=0%.
-$\mu=0.1,\lambda=0.04$.
-(a) Synthetic(iid); (b) Synthetic(0,0); (c) Synthetic(0.5,0.5); (d) Synthetic(1,1).
-  • Impact of $\alpha$
-    Figure 10: Impact of $\alpha$ on Femnist. Trained with fraction=1, straggler=0%. $\mu=0.1,\ \lambda=0.04$.
-H.5.2 Output perturbation
-Figure 11: Output perturbation on synthetic datasets. Trained with fraction=1, straggler=0%. $\mu=0.1,\lambda=0.04$. (a) Synthetic(iid); (b) Synthetic(0,0); (c) Synthetic(0.5,0.5); (d) Synthetic(1,1).
-Figure 12: Objective and output perturbation on Femnist and Sent140.
-(a) shows the result using objective perturbation on Femnist.
-The average accuracy of private Upcycled-FL is 71.04%(70.71%), of private FedProx is 19.72%(18.82%).
-(b) shows the result using output perturbation on Femnist.
-The average accuracy of private Upcycled-FL is 79.43%(79.63%), of private FedProx is 78.43%(78.36%).
-(c) shows the result using objective perturbation on Sent140.
-The average accuracy of private Upcycled-FL is 74.36%(72.38%), of private FedProx is 69.94%(68.02%).
-(d) shows the result using output perturbation on Sent140.
-The average accuracy of private Upcycled-FL is 74.42%(72.52%), of private FedProx is 62.10%(62.13%).All trained with fraction=1, straggler=0%, $\mu=0.1,\lambda=0.04$.
-For objective perturbation, $\delta=0.001$.
+F.2 Details of Algorithm FedProx
+Here we present the detailed algorithm of FedProx:
+Algorithm 2 FedProx [Li et al., 2020]
+1:  Input: $\mu>0$, $\{\mathcal{D}_{i}\}_{i\in\mathcal{I}}$, $\overline{\omega}^{0}$
+2:  for $t=1$ to $T$ do
+3:     The central server sends the current global model parameter $\overline{\omega}^{t}$ to all the clients.
+4:     A subset of clients get active and each active client updates its local model by finding (approximate) minimizer of local loss function:
+$\omega_{i}^{t+1}=\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{t}||^{2}.$
+5:     Each client sends its local model to server.
+6:     The central server updates the global model by aggregating all local models:
+$\overline{\omega}^{t+1}=\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t+1}.$
+7:  end for
+F.3 Convergence on all datasets
+Here we present the convergence and accuracy of the testing dataset in the final iteration for all datasets under 90% straggler and 30% straggler scenarios.
+F.3.1 90% Straggler
+(a) Syn(iid)
+(b) Syn(0,0)
+(c) Syn(0.5,0.5)
+(d) Syn(1,1)
+(e) FEMNIST
+(f) Sent140
+Figure 5: Convergence of Upcycled-FL and regular FL methods with the approximate same training time under 90% Straggler.
+F.3.2 30% Straggler
+We also conduct an experiment with a 30% straggler scenario to examine the convergence of our method.
+Table 3 demonstrates that our conclusions remain valid even when the straggler rate is relatively low.
+Table 3: Average accuracy with 30% straggler on the testing dataset over four runs, the experimental setting is same as Table 1.
+Method Dataset
+Syn(iid) Syn(0,0) Syn(0.5,0.5) Syn(1,1) FEMNIST Sent140
+FedAvg 98.50±0.18 80.30±1.01 82.60±0.72 78.81±2.10 83.06±1.30 74.29±0.12
+Upcycled-FedAvg 99.30±0.25 80.77±1.09 83.09±0.14 80.74±2.71 84.41±0.09 75.18±0.50
+FedAvgM 98.50±0.07 81.08±2.33 82.89±0.12 80.64±2.64 81.60±4.57 74.61±2.28
+Upcycled-FedAvgM 99.12±0.17 82.20±1.80 82.68±0.19 80.43±2.51 82.31±3.17 74.16±2.84
+FedProx 97.22±0.24 80.88±0.67 82.31±1.05 80.13±2.76 81.57±1.46 74.23±0.31
+Upcycled-FedProx 98.24±0.21 81.52±0.82 83.05±0.14 81.73±0.84 82.85±1.60 74.83±0.14
+Scaffold 98.28±0.14 79.95±1.67 79.54±0.72 72.27±4.68 76.85±5.55 75.76±0.21
+Upcycled-Scaffold 99.52±0.18 79.46±1.62 81.83±2.08 73.34±1.13 77.69±2.39 77.01±0.30
+FedDyn 98.17±0.08 82.06±0.61 80.97±1.04 78.65±3.83 84.39±0.95 75.88±0.11
+Upcycled-FedDyn 98.57±0.22 82.08±0.68 82.80±1.42 81.11±2.64 85.34±2.47 75.90±0.33
+pFedme 96.45±0.14 91.05±0.60 89.64±1.04 92.88±0.67 76.15±3.08 72.87±0.26
+Upcycled-pFedme 96.89±0.27 91.06±0.61 89.40±1.13 92.64±0.90 77.36±3.59 73.48±0.49
+FedYogi 99.38±0.28 81.06±2.53 80.65±1.32 79.14±1.63 79.98±7.83 77.63±0.29
+Upcycled-FedYogi 99.60±0.14 81.77±2.06 81.72±1.02 80.32±0.76 81.96±4.28 77.78±0.85
+F.3.3 30% Straggler
+(a) Syn(iid)
+(b) Syn(0,0)
+(c) Syn(0.5,0.5)
+(d) Syn(1,1)
+(e) FEMNIST
+(f) Sent140
+Figure 6: Convergence of Upcycled-FL and regular methods with the approximate same training time under 30% Straggler.
+F.4 Additional Privacy Experiments
+In addition to privacy experiments on synthetic datasets, we also report the output perturbation and objective perturbation on real-world datasets: FEMNIST and Sent140 in Figure 7 and 8.
+Due to limited computational resources, we conduct these experiments using a subset of the previously established baselines.
+Figure 7: Comparison of private Upcycled-FL and private FL methods using output perturbation (left) and objective perturbation (right) on FEMNIST.
+Figure 8: Comparison of private Upcycled-FL and private FL methods using output perturbation (left) and objective perturbation (right) on Sent140.
+F.5 Experiments on full-scale lowercase letter of FEMNIST
+We conduct a convergence experiment with full-scale lowercase letters (26 classes) from the FEMNIST dataset, as shown in Table 4.
+We also conduct privacy experiments comparing our method with FedProx, as shown in Figure 9.
+Table 4: Average accuracy on the large-scale FEMNIST dataset
+Method FEMNIST(lowercase letters)
+30% straggler 90% straggler
+FedAvg $90.90\pm 4.55$ $88.50\pm 5.07$
+Upcycled-FedAvg $91.26\pm 4.06$ $88.69\pm 4.94$
+FedAvgM $91.02\pm 4.47$ $89.13\pm 4.35$
+Upcycled-FedAvgM $91.05\pm 4.36$ $89.28\pm 4.05$
+FedProx $91.91\pm 0.38$ $89.02\pm 4.53$
+Upcycled-FedProx $92.08\pm 0.72$ $89.44\pm 4.30$
+Scaffold $89.69\pm 3.92$ $89.11\pm 3.74$
+Upcycled-Scaffold $89.77\pm 4.01$ $89.72\pm 3.90$
+FedDyn $92.88\pm 0.30$ $91.42\pm 0.24$
+Upcycled-FedDyn $93.56\pm 0.30$ $92.64\pm 0.37$
+pFedme $81.36\pm 4.63$ $80.72\pm 1.47$
+Upcycled-pFedme $85.62\pm 1.43$ $85.14\pm 0.93$
+FedYogi $87.64\pm 1.05$ $85.61\pm 0.56$
+Upcycled-FedYogi $88.89\pm 1.84$ $86.48\pm 1.30$
+Figure 9: Comparison of private Upcycled-FedProx and private FedProx using output perturbation (left) and objective perturbation (right).
+F.6 Training Time
+We report the average training time to compare the communication cost in Table 5.
+Table 5: Training time for output perturbation experiments on Syn(iid).
+Time/s FedAvg FedAvgM FedProx Scaffold FedDyn pFedme FedYogi
+Baseline $147.50\pm 6.69$ $147.47\pm 7.49$ $139.91\pm 10.81$ $153.88\pm 12.70$ $181.49\pm 17.60$ $467.42\pm 12.87$ $153.53\pm 13.72$
+Upcycled $86.44\pm 9.23$ $84.95\pm 3.39$ $80.59\pm 3.29$ $95.09\pm 13.89$ $99.82\pm 3.32$ $259.54\pm 18.27$ $92.82\pm 2.24$